### v1

# Unveiling Hallucination in Text, Image, Video, and Audio Foundation Models: A Comprehensive Review

Abstract
The rapid advancement of foundation models (FMs) across language, image, audio, and video domains has shown remarkable capabilities in diverse tasks. However, the proliferation of FMs brings forth a critical challenge: the potential to generate hallucinated outputs, particularly in high-stakes applications. The tendency of foundation models to produce hallucinated content arguably represents the biggest hindrance to their widespread adoption in real-world scenarios, especially in domains where reliability and accuracy are paramount. This survey paper presents a comprehensive overview of recent developments that aim to identify and mitigate the problem of hallucination in FMs, spanning text, image, video, and audio modalities. By synthesizing recent advancements in detecting and mitigating hallucination across various modalities, the paper aims to provide valuable insights for researchers, developers, and practitioners. Essentially, it establishes a clear framework encompassing definition, taxonomy, and detection strategies for addressing hallucination in multimodal foundation models, laying the foundation for future research in this pivotal area.
1 Introduction
The rapid progress in large-scale foundation models (FMs), spanning language, image, audio, and video domains, has revolutionized the field of artificial intelligence (AI). Models such as GPT-3 Brown et al. (2020), MiniGPT-4 Zhu et al. (2023), AudioLLM Borsos et al. (2023), and LaViLa Zhao et al. (2022) have demonstrated remarkable abilities across diverse tasks, from text generation to multimodal understanding. As these models find wider applications in critical domains, there is a growing imperative to comprehend and alleviate their propensity to produce hallucinated outputs.
1.1 Hallucination
Hallucination refers to instances where FMs generate content that appears plausible and emulates human-like patterns but lacks a coherent understanding of the underlying context or factual grounding Xu et al. (2024b). These hallucinated outputs can manifest in various forms, ranging from trivial factual inaccuracies to more severe cases where the model generates entirely imaginary or absurd content. This issue is not limited to textual modalities but extends across diverse domains, including images, videos, and audio generated by large-scale foundation models. The root causes of hallucination are multifaceted, potentially stemming from biases in training data, limited access to current information, or the model’s inherent constraints in comprehending and generating contextually precise responses. Deploying these powerful models without addressing their hallucination tendencies can have severe consequences, perpetuating misinformation, leading to incorrect conclusions, and potentially causing adverse effects in critical applications. Addressing hallucinations in FMs has become an active research area and a key focus for development efforts, with ongoing efforts to mitigate this behavior. Strategies employed include fine-tuning models on domain-specific data, leveraging diverse training data to enhance model robustness, and developing improved evaluation metrics to identify and reduce hallucination tendencies.
1.2 Types of Hallucination
Hallucinations in large foundation models can manifest in various forms, each posing unique challenges and requiring tailored strategies for mitigation. These types of hallucinations can be categorized as follows: Contextual disconnection Zhang et al. (2023d) describes a situation in which the output or content produced by a model across different modalities is inconsistent or out of sync with the context that the user or the input data have provided or expected. Semantic distortion  Tjio et al. (2022) refers to a type of inconsistency or error in generated content where the semantics or underlying meaning of the input is misrepresented or altered in the output. Content hallucination Moernaut et al. (2018)is the term used to describe a phenomenon seen in generative models when features or elements that are generated as output are either unreal given the context or absent from the input data. Factual inaccuracy  Zhang et al. (2023d) describes a kind of error seen in generative models when information that is inaccurate, deceptive, or at odds with the known facts appears in the generated output. Figure 1 illustrates various types of hallucinations.
Figure 1: Types of hallucinations illustrated: This image shows several types of hallucinations. Proper explanations of hallucinations are indicated as hallucinated elements (HE) and are highlighted in bold red text.
1.3 Motivation and Contributions
Most of the existing survey papers have explored hallucination in the context of large language models (LLMs) Huang et al. (2023)Tonmoy et al. (2024). Recent studies have shown that hallucination can also occur in vision, audio, and video foundation models, highlighting the need for a comprehensive understanding of this challenge across multiple modalities Liu et al. (2024a)Sahoo et al. (2024)Rawte et al. (2023b). To address this gap, our survey aims to provide a holistic and multimodal perspective on the hallucination challenge in FMs. This review comprehensively examines existing research across language, vision, video, and audio domains to understand the mechanisms, detection methods, and mitigation strategies for hallucination in FMs. It serves as a vital resource for researchers and practitioners, aiding in the development of more robust AI solutions. Additionally, it includes a detailed taxonomy diagram in Figure. 2 and a summarized Table 1 illustrating recent advancements across different modalities. The contributions of this survey paper are as follows:
  • Establish a precise definition and structured taxonomy of hallucination in the context of large-scale foundation models.
  • Identify the key factors and mechanisms that contribute to the emergence of hallucination across different modalities.
  • We have presented the various detection and mitigation strategies that have been proposed to address the hallucination problem in a multimodal setting.
  • We have provided a comprehensive summary of the methodologies pertaining to hallucination techniques in large foundational models in Table 1, detailing their approaches to hallucination detection, mitigation, task considerations, datasets utilized, and evaluation metrics employed. This will offer readers a concise overview of recent advancements in this field.
The paper is structured as follows: Section 2 delves into hallucinations in LLMs, while Section 3 explores large vision language models. Additionally, Section 4 and Section 5 examine hallucinations in large video and large audio models, respectively. Section 6 presents a discourse on the ramifications of hallucinations, exploring whether they pose advantages or disadvantages. Subsequently, Section 7 addresses the limitations of the current work, while Section 8 outlines potential future directions. Finally, Section 9 presents the concluding remarks.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=center, text centered, minimum width=5em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style= rotate=90, child anchor=north, parent anchor=south, anchor=center , , where level=1text width=15em,font=,, where level=2text width=14em,font=,, where level=3minimum width=10em,font=,, where level=4text width=26em,font=,, where level=5text width=20em,font=,, [ Hallucinations, for tree=fill=bg16, text width=14em [ Text §2, for tree=fill=red!50, text width=18em [ Detection, for tree=fill=bg1, text width=13.5em [ FACTOID Rawte et al. (2024b), for tree=fill=pg56,text width=28em ] [ FACTOR Muhlgay et al. (2023), for tree=fill=pg56,text width=28em ] [ FactCHD Chen et al. (2023b), for tree=fill=pg56,text width=28em ] [ HalluQA Cheng et al. (2023), for tree=fill=pg56,text width=28em ] [ HaluEval Li et al. (2023b), for tree=fill=pg56,text width=28em ] [ SelfCheckGPT Manakul et al. (2023), for tree=fill=pg56,text width=28em ] [ FACTSCORE Min et al. (2023), for tree=fill=pg56,text width=28em ] [ FacTool Chern et al. (2023), for tree=fill=pg56,text width=28em ] ] [ Mitigation, for tree=fill=bg2, text width=13.5em [ DoLa Chuang et al. (2023), for tree=fill=bg4,text width=28em ] [ CoK Li et al. (2023d), for tree=fill=bg4,text width=28em ] [ MixAlign Knowledge , for tree=fill=bg4,text width=28em ] [ SELF-FAMILIARITY Luo et al. (2023), for tree=fill=bg4,text width=28em ] [ CoVe Dhuliawala et al. (2023), for tree=fill=bg4,text width=28em ] [ HALOCHECK Elaraby et al. (2023), for tree=fill=bg4,text width=28em ] [ Instructional Prompting Varshney et al. (2023), for tree=fill=bg4,text width=28em ] [ PURR Chen et al. (2023a), for tree=fill=bg4,text width=28em ] [ LLM-AUGMENTER Peng et al. (2023), for tree=fill=bg4,text width=28em ] ] ] [ Image §3, for tree=fill=orange!50, text width=18em [ Detection, for tree=fill=bg1, text width=13.5em [ HallusionBench Guan et al. (2023), for tree=fill=dandyshade,text width=28em ] [ GAVIE Liu et al. (2023), for tree=fill=dandyshade,text width=28em ] [ PhD Liu et al. (2024b), for tree=fill=dandyshade,text width=28em ] [ VHTest Huang et al. (2024), for tree=fill=dandyshade,text width=28em ] [ M-HalDetect Gunjal et al. (2024), for tree=fill=dandyshade,text width=28em ] [ ChartBench Xu et al. (2023b), for tree=fill=dandyshade,text width=28em ] [ MSG-MCQ Lu et al. (2024), for tree=fill=dandyshade,text width=28em ] [ MMVP Tong et al. (2024), for tree=fill=dandyshade,text width=28em ] [ REVO-LION Liao et al. (2023), for tree=fill=dandyshade,text width=28em ] [ CIEM Hu et al. (2023), for tree=fill=dandyshade,text width=28em ] [ FAITHSCORE Jing et al. (2023), for tree=fill=dandyshade,text width=28em ] [ POPE Li et al. (2023f), for tree=fill=dandyshade,text width=28em ] [ VQA Changpinyo et al. (2022), for tree=fill=dandyshade,text width=28em ] [ HaELM Wang et al. (2023), for tree=fill=dandyshade,text width=28em ] [ NOPE Lovenia et al. (2023), for tree=fill=dandyshade,text width=28em ] [ MMHAL-BENCH Sun et al. (2023), for tree=fill=dandyshade,text width=28em ] [ TouchStone Bai et al. (2023), for tree=fill=dandyshade,text width=28em ] ] [ Mitigation, for tree=fill=bg2, text width=13.5em [ HalluciDoctor Yu et al. (2023a), for tree=fill=bg8,text width=28em ] [ LURE Zhou et al. (2023), for tree=fill=bg8,text width=28em ] [ MARINE Zhao et al. (2024), for tree=fill=bg8,text width=28em ] [ FDPO Gunjal et al. (2024), for tree=fill=bg8,text width=28em ] [ ViGoR Yan et al. (2024), for tree=fill=bg8,text width=28em ] [ HA-DPO Zhao et al. (2023), for tree=fill=bg8,text width=28em ] [ VIGC Wang et al. (2024), for tree=fill=bg8,text width=28em ] [ Data Centric Approach Lu et al. (2024), for tree=fill=bg8,text width=28em ] [ MoF Tong et al. (2024), for tree=fill=bg8,text width=28em ] [ InternLM-XComposer Zhang et al. (2023b), for tree=fill=bg8,text width=28em ] [ VCD Leng et al. (2023), for tree=fill=bg8,text width=28em ] [ Factually Augmented RLHF Sun et al. (2023), for tree=fill=bg8,text width=28em ] [ ObjMLM Dai et al. (2022), for tree=fill=bg8,text width=28em ] ] ] [ Video §4 , for tree=fill=bg5, text width=18em [ Detection, for tree=fill=bg1, text width=13.5em [ EMScore Shi et al. (2022), for tree=fill=bg7,text width=28em ] ] [ Mitigation, for tree=fill=bg2, text width=13.5em [ FactVC Liu and Wan (2023), for tree=fill=bg13,text width=28em ] [ CLearViD Chuang and Fazli (2023), for tree=fill=bg13,text width=28em ] [ MGAT He et al. (2022), for tree=fill=bg13,text width=28em ] ] ] [ Audio §5 , for tree=fill=lime!50, text width=18em [ Detection, for tree=fill=bg1, text width=13.5em [ PAM Deshmukh et al. (2024), for tree=fill=bg15,text width=28em ] [ CompA Ghosh et al. (2023), for tree=fill=bg15,text width=28em ] ] [ Mitigation, for tree=fill=bg2, text width=13.5em [ MusicLDM Chen et al. (2024), for tree=fill=bg14,text width=28em ] [ Re-AudioLDM Yuan et al. (2024), for tree=fill=bg14,text width=28em ] [ SECap Xu et al. (2024a), for tree=fill=bg14,text width=28em ] [ RECAP Ghosh et al. (2024), for tree=fill=bg14,text width=28em ] [ Cacophony Zhu and Duan (2024), for tree=fill=bg14,text width=28em ] [ EnCLAP Kim et al. (2024), for tree=fill=bg14,text width=28em ] ] ] ]
Figure 2: Taxonomy of hallucination in large foundation models, organized around detection and mitigation techniques.
Figure 3: LLM responses showing the types of hallucinations, highlighted in red, green, and blue Zhang et al. (2023d).
2 Hallucination in Large Language Models
Despite the progress of LLMs, a notable challenge persists in their proneness to hallucinate, impeding their practical implementation. For instance, the illustration in Figure 3 exemplifies the generated response by the LLM, showcasing indications of hallucination.
2.1 Hallucination Detection and Mitigation
Identifying hallucinations in LLMs is crucial for ensuring the credibility and reliability of their results, especially in scenarios requiring factual correctness. Existing fact-checking methods often rely on complex modules or external databases, requiring either output probability distributions or interfacing with external sources. SelfCheckGPT Manakul et al. (2023) offers a zero-resource black-box solution for detecting hallucinations in any LLM without relying on external resources. This method operates on the principle that an LLM familiar with a topic will produce consistent and comparable facts in its responses. In contrast, randomly sampled responses from an unfamiliar topic are likely to contain contradicting and hallucinated facts. Continuing the exploration of methods for passage-level hallucination detection, Yang et al. (2023) proposed a novel self-check approach based on reverse validation, aiming to identify factual errors without external resources automatically. They introduced a benchmark, Passage-level Hallucination Detection (PHD), generated using ChatGPT and annotated by human experts to assess different methods. Assessing the accuracy of long text generated by LLMs is challenging because it often contains both accurate and inaccurate information, making simple quality judgments insufficient. To address this, Min et al. (2023) introduced FACTSCORE (Factual Precision in Atomicity Score), a novel evaluation method that breaks down text into individual facts and measures their reliability. Huang and Chang (2023) introduced a unique strategy to mitigate hallucination risks in LLMs by drawing parallels with established web systems. They identified the absence of a "citation" mechanism, which refers to acknowledging or referencing sources or evidence, as a significant gap.
Addressing the need to identify factual inaccuracies in LLM-generated content, Rawte et al. (2024b) develop a multi-task learning (MTL) framework, integrating advanced long text embeddings like e5-mistral-7b-instruct, along with models such as GPT-3, SpanBERT, and RoFormer. This MTL approach exhibited a substantial performance gain, achieving an average accuracy improvement of 40% on the FACTOID benchmark when compared to leading textual entailment methods. Hallucination mitigation efforts have predominantly relied on empirical methods, leaving uncertainty regarding the possibility of complete elimination. To tackle this challenge, Xu et al. (2024b) introduces a formal framework defining hallucination as discrepancies observed between computable LLMs and a ground truth function. Through this framework, the study examines existing hallucination mitigation strategies and their practical implications for real-world LLM deployment. Rawte et al. (2024c) introduced the Sorry, Come Again (SCA) prompting technique to address hallucination in contemporary LLMs. SCA enhances comprehension by employing optimal paraphrasing and injecting [PAUSE] tokens to delay LLM generation. It analyzes linguistic nuances in prompts and their impact on the hallucinated generation, highlighting the difficulties posed by prompts characterized by lower readability, formality, or concreteness. Rawte et al. (2023a) investigate how LLMs respond to factually correct and incorrect prompts, categorizing their hallucinations into mild, moderate, and alarming subcategories. Additionally, the paper introduced the Hallucination eLiciTation dataset, comprising 75,000 text snippets annotated by humans, and introduced a novel Hallucination Vulnerability Index metric.
2.2 Domain Specific Works
Hallucinations pose severe risks in critical fields such as healthcare, finance, and law. Reliability and accuracy are crucial in these sectors, as any form of hallucination can lead to significant and adverse consequences.
2.2.1 Healthcare
In response to the hallucinations in the medical domain LLMs, Pal et al. (2023) introduced the Medical Domain Hallucination Test (Med-HALT), a specialized benchmark dataset aimed at assessing and mitigating hallucinations. Med-HALT consists of a varied multinational dataset drawn from medical records spanning numerous countries, encompassing a total of seven datasets. Ahmad et al. (2023) outlined essential steps for creating dependable, trustworthy, and unbiased models, emphasizing the need for quantifying, validating, and mitigating hallucinations within the healthcare context. Ji et al. (2023) introduced an interactive self-reflection approach aimed at improving the accuracy and coherence of answers generated by medical question-answering systems using LLMs. Through knowledge acquisition and feedback on answer generation, this methodology enhances the factuality, consistency, and logical progression of responses.
2.2.2 Finance
An empirical investigation explored the inclination of LLMs to generate hallucinations while engaging in financial activities.  Kang and Liu (2023) conducted an empirical investigation into the hallucination tendencies of LLMs in financial tasks. Their study assessed LLMs’ proficiency in explaining financial concepts, querying historical stock prices, and examined the efficacy of methods like few-shot learning and prompt-based tool learning in mitigating hallucination.  Roychowdhury et al. (2023) proposed a novel Langchain-based approach aimed at transforming data tables into hierarchical textual data chunks, facilitating versatile financial question answering. The framework involves classifying user queries by intention, retrieving relevant data chunks, generating customized LLM prompts, and evaluating responses for hallucinations and confidence.
2.2.3 Legal
The conventional approach to abstractive text summarization typically employs an encoder-decoder architecture, wherein the encoder encapsulates the essence of the source text while the decoder generates the summary. However, this method may produce summaries containing irrelevant or inaccurate information, which poses a significant concern in legal contexts where accuracy is crucial. To address these issues, Feijo and Moreira (2023) introduced LegalSumm, which creates distinct "views" of the source text, trains summarization models to produce independent summaries, and employs an entailment module to assess their fidelity to the source. Deroy et al. (2023) investigated the readiness of LLMs for generating abstractive summaries of case judgments by applying SOTA models to Indian court cases. While abstractive models generally scored slightly higher, the authors noted inconsistencies and hallucinations in the generated summaries. Understanding the meaning of open-ended legal terms is very important for legal professionals. They often look at how such terms were used and interpreted in previous court cases.  Savelka et al. (2023) evaluate GPT-4’s performance in generating factually accurate, clear, and relevant explanations of legal terms in legislation. A comparison is made between a baseline approach, where GPT-4 directly explains a legal term, and an augmented approach employing a legal information retrieval module to provide contextual sentences from case law.  Dahl et al. (2024) presents the initial evidence on the frequency and types of inaccuracies in the legal domain, providing valuable insights for evaluating LLMs in legal contexts. By examining the structured format of American case law, the study assesses three major LLMs: GPT-3.5, PaLM 2, and Llama.
2.3 Benchmark Evaluation
In certain instances, LLMs engage in a phenomenon termed "hallucination snowballing," where they fabricate false claims to rationalize prior hallucinations despite acknowledging their inaccuracy. To empirically explore this phenomenon, Zhang et al. (2023a) devise three question-answering datasets spanning diverse domains, wherein ChatGPT and GPT-4 often furnish inaccurate answers alongside explanations featuring at least one false claim. Significantly, the study suggested that the language model can discern these false claims as incorrect. Another benchmark dataset, FactCHD Chen et al. (2023b), was introduced to detect fact-conflicting hallucinations within intricate inferential contexts. It encompasses a range of datasets capturing different factual patterns and integrates fact-based evidence chains to improve assessment accuracy. Li et al. (2023b) introduced a dataset to assess the capability of LLMs to identify and recognize hallucinated or incorrect information. The outcomes highlighted ChatGPT’s inclination to produce hallucinated content, particularly on certain topics, introducing unverifiable information.
Figure 4: Example of types of Vision Hallucination.  Liu et al. (2024b) including visuals and the matching question-answer pairs and hallucination elements. While red annotated words do not exist or do not correspond to elements within the image, whereas words highlighted in green have correspondences within the image. Question and Answer are denoted by the letters Q and A, respectively.
3 Hallucination in Large Vision-Language Models
Large Vision-Language Models (LVLMs) have garnered significant attention in the AI community for their capacity to handle visual and textual data simultaneously. Nonetheless, similar to LLMs, LVLMs also confront the issue of hallucination. Figure 4 illustrates an example of visual hallucination.
3.1 Hallucination Detection and Mitigation
Dai et al. (2022) investigate the issue of object hallucinations in Vision-Language Pre-training (VLP) models, where textual descriptions generated by these models contain non-existent or inaccurate objects based on input images.  Li et al. (2023f) reveal widespread and severe object hallucination issues and suggest that visual instructions may influence hallucination. They observe that objects frequently depicted in visual instructions or co-occurring with image objects are more susceptible to hallucination. To enhance the evaluation process of object hallucination, the authors introduced a polling-based query method called POPE, which demonstrates improved stability and flexibility in assessing object hallucination. The absence of a standardized metric for assessing object hallucination has hindered progress in understanding and addressing this issue. To address this gap, Lovenia et al. (2023) introduced NOPE (Negative Object Presence Evaluation), a benchmark for evaluating object hallucination in vision-language models (VLMs) through visual question answering (VQA). Utilizing LLMs, the study generated a dataset of 29.5k synthetic negative pronoun (NegP) instances for NOPE. It thoroughly assessed the capability of 10 VLMs in detecting the absence of objects in visual questions, in addition to their typical performance on visual questions across nine other VQA datasets. Existing research focuses primarily on object hallucination, overlooking the LVLM hallucinations. Liu et al. (2024b) delve into Intrinsic Vision-Language Hallucination (IVL-Hallu) and proposes several novel IVL-Hallu tasks, including attribute, object, multi-modal conflicting, and counter-common-sense hallucination. They introduced a challenging benchmark dataset to assess and explore IVL-Hallu, conducting experiments on five LVLMs that revealed their limited effectiveness in addressing the proposed tasks. To mitigate object hallucination in LVLMs without resorting to costly training or API reliance, Zhao et al. (2024) introduced MARINE, a training-free and API-free solution. MARINE enhances LVLMs’ visual comprehension by combining existing open-source vision models and leveraging guidance without classifiers to incorporate object grounding features, thereby enhancing the precision of generated outputs. Evaluations across six LVLMs revealed MARINE’s effectiveness in reducing hallucinations and enhancing output detail, validated through assessments using GPT-4V.
HalluciDoctor Yu et al. (2023a) tackled hallucinations in Multi-modal Large Language Models (MLLMs) by using human error detection to identify and eliminate various types of hallucinations. Through rebalancing data distribution via counterfactual visual instruction expansion, they successfully mitigate 44.6% of hallucinations while maintaining competitive performance. Despite proficiency in visual semantic comprehension and meme humor, MLLMs struggle with chart analysis and understanding. Addressing this, Xu et al. (2023b) propose ChartBench, a benchmark assessing chart comprehension. ChartBench exposes MLLMs’ limited reasoning with complex charts, prompting the need for novel evaluation metrics like Acc+ and a handcrafted prompt, ChartCoT. Zhang et al. (2023b) introduced InternLM-XComposer, an LVLM aimed at designed to address the challenge of hallucination in image-text comprehension and composition. The performance of InternLM-XComposer’s text-image composition is evaluated through a robust procedure involving both human assessment and comparison to GPT4-Vision, with the model demonstrating competitive performance against solutions like GPT4-V and GPT3.5.
Cutting-edge LVLMs like InstructBLIP Dai et al. (2024), while generating visually grounded responses, often include inaccuracies like fictional items and flawed relationships. To enhance accuracy, Gunjal et al. (2024) introduced M-HalDetect, a pioneering multi-modal fine-grained hallucination detection dataset. This dataset serves as a benchmark for training LVLMs, leading to more precise outputs. Using fine-grained multi-modal reward models and enhancing FDPO significantly reduced InstructBLIP’s hallucination rate. These methods not only improve accuracy across LVLMs like LLaVA and mPLUG-OWL but also highlight the versatility and effectiveness of M-HalDetect in identifying and minimizing hallucinations.
Despite advancements in multi-modal tasks, LMMs often generate descriptions that are inconsistent with the accompanying image or human instructions. To address this, Liu et al. (2023) develop LRV-Instruction, a comprehensive dataset comprising 400k visual instructions across 16 tasks. This dataset includes both positive and negative instructions in various styles and semantic levels. Through LRV-Instruction, the issue of hallucination in existing LMMs was extensively examined, confirming its effectiveness in enhancing visual instruction tuning. Additionally, they introduced GAVIE, a novel method for evaluating visual instruction tuning without the need for human-labeled answers, which can be adapted to different types of instructions. The LVLM Hallucination Revisor (LURE) algorithm Zhou et al. (2023) was developed to correct object hallucination in LVLMs by refining the descriptions to produce more accurate and less hallucinatory outputs. Its methodology is based on an in-depth statistical analysis that identifies key factors contributing to object hallucination, such as the co-occurrence of certain objects in images, the uncertainty associated with objects during LVLM decoding, and the tendency for hallucinations to occur towards the end of the generated text. LURE is designed for seamless integration with a variety of LVLMs. When tested across multiple LVLMs, the integration of LURE resulted in significant enhancements in object hallucination correction, consistently outperforming other methods in both GPT and human evaluations based on various metrics.
3.2 Benchmark Evaluation
The current methods of developing LVLMs rely heavily on annotated benchmark datasets, which can exhibit domain bias and limit model generative capabilities. To address this, Li et al. (2023e) propose a novel data collection approach that synthesizes images and dialogues synchronously for visual instruction tuning, yielding a large dataset of image-dialogue pairs and multi-image instances.  Huang et al. (2024) introduced VHTest, a benchmark dataset with 1,200 diverse visual hallucinations (VH) instances across 8 VH modes. Evaluation of three SOTA MLLMs showed varying performance, with GPT-4V exhibiting lower hallucination than MiniGPT-v2. Rawte et al. (2024a) categorized visual hallucination in VLMs into eight orientations and introduced a dataset of 2,000 samples covering these types. They proposed three main categories of methods to mitigate hallucination: data-driven approaches, training adjustments, and post-processing techniques. Additionally, Wang et al. (2024) proposed the Visual Instruction Generation and Correction (VIGC) framework to address the shortage of high-quality instruction-tuning data for MLLMs. VIGC enables MLLMs to generate diverse instruction-tuning data while iteratively refining its quality through Visual Instruction Correction (VIC), mitigating hallucination risks. The framework produces diverse, high-quality data for fine-tuning models, validated through evaluations, improving benchmark performance, and overcoming language-only data limitations.
Figure 5: A video featuring descriptions generated by VLTinT model and ground truth (GT) with description errors highlighted in red italics Chuang and Fazli (2023).
4 Hallucinations in Large Video Models
Large Video Models (LVMs) represent a significant advancement, allowing for the processing of video data at scale. Despite their potential for various applications like video understanding and generation, LVMs face challenges with hallucinations, where misinterpretations of video frames can result in artificial or inaccurate visual data. This issue arises due to the video data complexity, which requires the model to process and comprehend it thoroughly. Figure 5 demonstrates the instances of hallucination observed in LVMs.
4.1 Hallucination Detection and Mitigation
The intricate task of dense video captioning, involving the creation of descriptions for multiple events within a continuous video, necessitates a thorough understanding of video content and contextual reasoning to ensure accurate description generation. However, this endeavor faces numerous challenges, potentially resulting in instances of inaccuracies and hallucinations Iashin and Rahtu (2020)Suin and Rajagopalan (2020). Traditional methods detect event proposals first, then caption subsets, risking hallucinations due to overlooking temporal dependencies. To address this, Mun et al. (2019) introduces a novel approach to modeling temporal dependencies and leveraging context for coherent storytelling. By integrating an event sequence generation network and a sequential video captioning network trained with reinforcement learning and two-level rewards, the model captures contextual information more effectively, yielding coherent and accurate captions while minimizing the risk of hallucinations.  Liu and Wan (2023) introduces a novel weakly-supervised, model-based factuality metric called FactVC, which outperforms previous metrics. Furthermore, they provide two annotated datasets to promote further research in assessing the factuality of video captions. Wu and Gao (2023) proposed a context-aware model that incorporates information from past and future events to influence the description of the current event conditionally. Their approach utilizes a robust pre-trained context encoder to encode information about the surrounding context events, which is then integrated into the captioning module using a gate-attention mechanism. Experimental findings on the YouCookII and ActivityNet datasets demonstrate that the proposed context-aware model outperforms existing context-aware and pre-trained models significantly. To enhance dense video captioning,  Zhou et al. (2024) introduces a streaming model comprising a memory module for long video handling and a streaming decoding algorithm enabling predictions before video completion. This approach notably boosts performance on prominent dense video captioning benchmarks, such as YouCook2, ActivityNet, and ViTT.
Video infilling and prediction tasks are crucial for assessing a model’s ability to comprehend and anticipate the temporal dynamics within video sequences Höppe et al. (2022). To address this, Himakunthala et al. (2023) introduced an inference-time challenge dataset containing keyframes with dense captions and structured scene descriptions. This dataset contains keyframes supplemented with unstructured dense captions and structured FAMOUS: (Focus, Action, Mood, Objects, and Setting) scene descriptions, providing valuable contextual information to support the models’ understanding of the video content. They employed various language models like GPT-3, GPT-4, and Vicuna with greedy decoding to mitigate hallucination risks.
Prominent developments in video inpainting have been observed recently, especially in situations where explicit guidance like optical flow helps to propagate missing pixels across frames Ouyang et al. (2021). However, difficulties and constraints occur from a lack of cross-frame information. Yu et al. (2023b) aims to tackle the opposite issue rather than depending on using pixels from other frames. The suggested method presents a Deficiency-aware Masked Transformer (DMT), a dual-modality-compatible inpainting framework. This approach improves handling scenarios with incomplete information by pre-training an image inpainting model to serve as a prior for training the video model.
Understanding scene affordances, which involve potential actions and interactions within a scene, is crucial for comprehending images and videos. Kulal et al. (2023) introduces a method for realistically inserting people into scenes. The model seamlessly integrates individuals into scenes by deducing realistic poses based on the context and ensuring visually pleasing compositions. Chuang and Fazli (2023) introduces CLearViD, a transformer-based model that utilizes curriculum learning techniques to enhance performance. By adopting this approach, the model acquires more robust and generalizable features. Furthermore, CLearViD incorporates the Mish activation function to address issues like vanishing gradients, thereby reducing the risk of hallucinations by introducing nonlinearity and non-monotonicity. Extensive experiments and ablation studies validate the effectiveness of CLearViD, with evaluations on ActivityNet Captions and YouCook2 datasets showcasing significant improvements over existing SOTA models in terms of diversity metrics.
4.2 Benchmark Evaluation
Zhang et al. (2006) created an innovative two-level hierarchical fusion method to hallucinate facial expression sequences from training video samples using only one frontal face image with a neutral expression. To effectively train the system, they introduced a dataset specifically designed for facial expression hallucination, which included 112 video sequences covering four types of facial expressions (happy, angry, surprise, and fear) from 28 individuals, resulting in the generation of reasonable facial expression sequences in both the temporal and spatial domains with less artifact. In the realm of video understanding, the development of end-to-end chat-centric systems has become a growing area of interest. Zhou et al. (2018) assembled the YouCook2 dataset, an extensive set of cooking videos with temporally localized and described procedural segments, to facilitate procedure learning tasks. Li et al. (2023c) introduced "VideoChat", a novel approach integrating video foundation models and LLMs through a learnable neural interface to enhance spatiotemporal reasoning, event localization, and causal relationship inference in video understanding. The researchers constructed a video-centric instruction dataset with detailed descriptions and conversations, emphasizing spatiotemporal reasoning and causal relationships. To counteract model hallucination, they employed a multi-step process to condense video descriptions into coherent narratives using GPT-4 and refined them to improve clarity and coherence. To explore the challenge of deducing scene affordances, Kulal et al. (2023) curated a dataset of 2.4M video clips, showcasing a variety of plausible poses that align with the scene context.
Figure 6: Audio hallucination examples for each classes - Type A: Involving hallucinations of both objects and actions $|$ Type B: Featuring accurate objects but hallucinated actions $|$ Type C: Displaying correct actions but hallucinated objects Nishimura et al. (2024).
5 Hallucinations in Large Audio Models
Large audio models (LAMs) have emerged as a powerful tool in the realm of audio processing and generation, with a wide range of applications like speech recognition, music analysis, audio synthesis, and captioning Latif et al. (2023)Ghosal et al. (2023). Although these models have demonstrated remarkable capabilities across various domains, they are susceptible to hallucinations. These anomalies can take several forms, from creating unrealistic audio by piecing together fabricated snippets to injecting false information, such as quotes or facts, into summaries. Additionally, they may fail to accurately capture the inherent features of audio signals, such as timbre, pitch, or background noise Shen et al. (2023).
5.1 Hallucination Detection and Mitigation
In the realm of audio captioning, where natural language descriptions for audio clips are automatically generated, a significant challenge arises from the over-reliance on the visual modality during the pre-training of audio-text models. This reliance introduced data noise and hallucinations, ultimately undermining the accuracy of the resulting captions. To address this issue, Xu et al. (2023a) introduces an AudioSet tag-guided model designed to bootstrap large-scale audio-text data (BLAT). Notably, this model sidesteps the incorporation of video, thus minimizing noise associated with the visual modality. The experimental findings across a range of tasks, including retrieval, generation, and classification, validate the effectiveness of BLAT in mitigating hallucination issues.
Speech emotions play a crucial role in human communication and find extensive applications in areas such as speech synthesis and natural language understanding. However, traditional categorization approaches may fall short of capturing the nuanced and intricate nature of emotions conveyed in human speech Jiang et al. (2019)Han et al. (2021)Ye et al. (2021). SECap Xu et al. (2024a), a framework designed for speech emotion captioning, aims to capture the intricate emotional nuances of speech using natural language. SECap utilizes various components, including LLaMA as the text decoder, HuBERT as the audio encoder, and Q-Former as the Bridge-Net, to generate coherent emotion captions based on speech features. Audio-language models, despite their capability for zero-shot inference, confront challenges like hallucinating task-specific details despite strong performance. To address this, Elizalde et al. (2024) introduces the Contrastive Language-Audio Pretraining (CLAP) model. Pre-trained with 4.6 million diverse audio-text pairs, CLAP features a dual-encoder architecture, enhancing representation learning for improved task generalization across sound, music, and speech domains.
5.2 Benchmark Evaluation
To address the scarcity of data in the specific domain of music captioning, Doh et al. (2023) introduced LP-MusicCaps, a comprehensive dataset comprising 0.5 million audio clips accompanied by approximately 2.2 million captions. Leveraging LLMs, they train a transformer-based music captioning model with the dataset and assess its performance under zero-shot and transfer-learning scenarios, demonstrating its superiority over supervised baseline models.  Nishimura et al. (2024) investigated audio hallucinations in large audio-video language models, where audio descriptions are generated primarily based on visual information, neglecting audio content. They have classified these hallucinations into three distinct types such as Involving hallucinations of both objects and actions, Featuring accurate objects but hallucinated actions, and Displaying correct actions but hallucinated objects as represented in Fig. 6. In their investigation, they gathered 1000 sentences by soliciting audio information and then annotated them to determine whether they contained auditory hallucinations, further categorizing the type of hallucination if detected. To assess compositional reasoning in LAMs, Ghosh et al. (2023) introduced CompA, consisting of two expert-annotated benchmarks primarily focused on real-world audio samples. This benchmark is employed to fine-tune CompA-CLAP with a novel learning approach, enhancing its compositional reasoning skills and demonstrating substantial improvement over all the baseline models in tasks requiring compositional reasoning.
TEXT Paper Detection Mitigation Task Dataset(s) Evaluation Metric(s)
Manakul et al. (2023) Yes No QA Wikibio Entropy
Li et al. (2022) Yes Yes QA, Dialog summarization Halueval Automatic
Mündler et al. Yes Yes Text generation Manual F1 Score
Chen et al. (2023a) No Yes Editing for attribution MCQ, Dialog Attribution, Preservation
Zhang et al. (2023c) No Yes Question knowledge alignment Fuzzy QA Attributable to Identified Sources
Zhang et al. (2023a) Yes No QA Manual Accuracy
Peng et al. (2023) No Yes Task-oriented dialog News, Customer service F1 Score, Bleu-4
Cui et al. (2023) No Yes QA Manual Ranking
Azaria and Mitchell (2023) Yes No Classification Manual Accuracy
Li et al. (2023d) Yes Yes Knowledge-intensive tasks Fever, QA Accuracy
Elaraby et al. (2023) Yes Yes Consistency, Actuality, QA Manual NBA domain Pearson Correlation Coefficient
Varshney et al. Yes Yes Text generation Wikibio Percentage of mitigated hallucination
Jha et al. (2023) Yes No Dialog N/A N/A
Pal et al. (2023) No No Reasoning hallucination Med-Halt Accuracy, Pointwise Score
McKenna et al. (2023) Yes No Textual entailment Altered Directional Inference Entailment Probability
Guerreiro et al. (2023) Yes Yes MT FLores 101, WMT ,TICO BLEU
Huang and Chang (2023) Yes Yes N/A N/A N/A
Luo et al. (2023) Yes Yes Concept extraction Concept-7 AUC, Accuracy, F1 Score
Gao et al. (2022) Yes Yes Editing attribution NQ, SQA Auto-AIS (Attr_auto)
Yang et al. (2023) Yes No
Detect factual
errors automatically
PHD, WikiBio-GPT3
Precision, Recall,
F1 Score, Accuracy
Min et al. (2023) Yes Yes Fact verification Manual(Wikipedia) FActScore
Rawte et al. (2024b) Yes Yes Factual inaccuracies detection FACTOID HV I_auto
Ahmad et al. (2023) Yes Yes Hallucination in healthcare N/A FActScores
Ji et al. (2023) Yes Yes Generative and knowledge-intensive
PubMedQA, MEDIQA2019,
MedQuAD, and MASH-QA
Unigram F1, ROUGE-L,
Med-NLI, and CTRLEval
Kang and Liu (2023) Yes Yes Hallucination in finance N/A FActScores
Roychowdhury (2024) No Yes QA N/A N/A
Savelka et al. (2023) No Yes Factual evaluation in legislation N/A N/A
Dahl et al. (2024) Yes No Legal hallucination Manual N/A
Rawte et al. (2024c) No Yes Comprehension enhancement SCA-90K Cosine similarity
IMAGE Li et al. (2023f) Yes No Evaluation of object hallucination MSCOCO CHAIR, POPE
Gunjal et al. (2024) Yes Yes VQA M-Hall Detect Accuracy
Dai et al. (2022) No Yes Image captioning CHAIR CIDEr
Lovenia et al. (2023) Yes No Object hallucination NOPE
METEOR, Exact match accuracy,
NegP Accuracy
Liu et al. (2024b) Yes No Intrinsic vision-language hallucination PhD Accuracy
Zhao et al. (2024) Yes Yes Non-existing object hallucination MSCOCO CHAIR, POPE, GPT-4V, recall
Huang et al. (2024) Yes No Visual hallucination YNQ, OEQ Accuracy
Rawte et al. (2024a) Yes No Video captioning ActivityNet-Fact, YouCook2-Fact FactVC
Wang et al. (2024) No Yes
Generate instruction data
for vision-language
VIGC-LLaVA-COCO,
VIGC-LLaVA-Objects365
Conv, Detail, Complex
Yu et al. (2023a) Yes Yes
Machine-generated
visual instruction
LLaVA-Instruction-158K CHAIR
Guan et al. (2023) No Yes Visual questions HallusionBench Accuracy
Liu et al. (2023) Yes Yes Vision language LRV-Instruction GAVIE
Xu et al. (2023b) Yes No
Evaluation of MLLMs on
chart comprehension
ChartBench Acc+
Lu et al. (2024) Yes Yes Vision language MSG-MCQ Accuracy
Tong et al. (2024) Yes No Visual question answering MMVP, VQA Accuracy
Liao et al. (2023) Yes No Vision language REVO-LION Meta Quality (MQ)
Hu et al. (2023) Yes Yes
Visual captioning,
Visual question answering
CIEM
Accuracy, Precision,
Recall, F1 Score
Jing et al. (2023) Yes No Meta-evaluation LLaVA-1k, MSCOCO-Cap FAITHSCORE
Changpinyo et al. (2022) No Yes Multilingual visual question answering MaXM Accuracy
Wang et al. (2023) Yes No Content generation N/A Precision, Recall, F1 Score
Sun et al. (2023) No Yes Visual-language alignment MMHAL-BENCH N/A
Bai et al. (2023) Yes No
Evaluate hallucination of
vision language model
TouchStone Hallucination Score
Zhou et al. (2023) No Yes Hallucination mitigation in LVMs MSCOCO CHAIR, BLEU, CLIP
Yan et al. (2024) No Yes Visual grounding MMViG
HL, CA, AA,
RA, RL, RS, DL
Zhao et al. (2023) Yes Yes Overcome hallucination in LVMs POPE, SHR Accuracy, Precision, F1 Score
Zhang et al. (2023b) No Yes
Image text comprehension
and composition
MMBench, SeedBench, QBench,
MMBench-CN, Chinese Bench
LR, AR, $R$,
FP-C, FP-S, CP
VIDEO Kulal et al. (2023) No Yes Affordance prediction Manual FID, FCKh
Himakunthala et al. (2023) No Yes Video infilling, Scene prediction Manual N/A
Li et al. (2023c) No Yes Visual dialogue Manual N/A
Zhou et al. (2024) No Yes Video captioning
ActivityNet Captions,
YouCook2, ViTT
CIDER, METEOR, SODAc
Höppe et al. (2022) Yes No Video prediction BAIR, Kinetics 600, UCF-101 Frechet Video Distance
Chuang and Fazli (2023) No Yes Video description Activity Net Captions, YouCook2
METEOR, ROUGE_L, CIDER,
BLEU_4, DIV-2, RE_4
AUDIO Li et al. (2023a) No Yes Classification Manual Mean avg precision
Doh et al. (2023) No Yes Audio captioning LP MusicCaps BLEU
Xu et al. (2023a) No Yes Caption generation AudioCaps R@K, COCO & FENCE
Liu and Wan (2023) No Yes Audio captioning MusciCaps BLEU
Nishimura et al. (2024) Yes No Evaluation of LAMs LAION_CLAP,MS_CLAP Recall, Precision, F1 Score
Table 1: Overview of the hallucination detection and mitigation landscape in FMs across modalities (Text, Image, Video, and Audio). Each work is categorized based on factors such as detection, mitigation, tasks, datasets, and evaluation metrics.
6 Hallucination: Good or Bad?
Hallucinations in large-scale models present a complex interplay between creativity and uncertainty. On one hand, the ability to traverse beyond conventional data boundaries can lead to the generation of novel and innovative outputs. Hallucinations can spark exploratory learning, revealing unexpected patterns and features within the data. They can also serve as a form of stress testing, improving the model’s robustness and adaptability. Furthermore, these unexpected outputs can even inspire human creativity, serving as a springboard for new ideas and perspectives Rawte et al. (2023b). However, this dual nature of hallucinations also introduced significant drawbacks. The quality and coherence of hallucinatory outputs can be questionable, posing challenges in applications where accuracy and reliability are paramount. Hallucinations can also propagate misinformation and biases present in the model’s training data, potentially reinforcing existing prejudices and eroding user trust. The reduced interpretability of these outputs can further undermine the model’s credibility and adoption. Ethical concerns arise when hallucinations produce inappropriate, offensive, or harmful content. Careful monitoring and control mechanisms are essential to prevent the generation of outputs that could cause harm or distress to users. Navigating this intricate balance between exploration and fidelity is crucial for maximizing the utility of large models while mitigating the risks associated with unexpected outputs. Overall, the phenomenon of hallucinations in large-scale models highlights the need for a nuanced understanding and strategic management of these capabilities.
7 Limitations
Previous survey papers primarily focused on hallucination in Large Language Models and did not extensively cover hallucinations in vision, audio, and video modalities. In this survey paper, our aim is to provide a comprehensive overview of hallucinations across all modalities, considering that hallucinations can occur in any large foundation model. Despite our efforts to provide a comprehensive summary of recent advancements related to hallucination techniques in all foundational models, we acknowledge that we may miss some relevant work in the field.
8 Future Directions
Researchers are actively investigating hallucination mitigation techniques as the challenge of hallucination can be crucial in sensitive areas where generating fictional or incorrect content could have serious consequences Tonmoy et al. (2024)Rawte et al. (2023b). Here are the potential directions for addressing this critical issue of hallucination in these FMs:
Data Resources: Recent studies have highlighted the efficacy of simple fine-tuning on carefully curated high-quality samples for reducing hallucinations, surpassing the impact of large-scale fine-tuning and reinforcement learning approaches. For knowledge-intensive domains, the development of entity-centered fine-tuned instructions that integrate structured knowledge from knowledge graphs shows promise in enhancing accuracy and relevance. Additionally, employing alignment techniques tailored to specific tasks or fields has proven effective in mitigating hallucinations. As research in this area progresses, more resources focused on improving alignment through task-specific or domain-adapted approaches are expected, further bolstering the reliability of language models in generating factual and trustworthy content.
Automated Evaluation: Developing specialized evaluation metrics that consider factors such as factual accuracy and coherence can be useful for hallucination detection. Combining automated evaluation with human judgments via crowdsourcing can capture nuanced aspects challenging for automated systems alone. Additionally, adversarial testing methodologies are also being developed to expose AI systems to crafted inputs, aiding in identifying weaknesses and enhancing resilience against hallucination. Furthermore, fine-tuning FMs on datasets emphasizing fact-checking and accuracy offers another avenue to improve content reliability and reduce the occurrence of hallucinations.
Improving Detection and Mitigation techniques: Mitigating hallucinations in FMs necessitates a multifaceted approach that leverages reasoning mechanisms, knowledge graph integration, specialized fact-checking models, bias mitigation techniques, and active learning methodologies. Emerging techniques like Chain of Thought (CoT) Wei et al. (2022) and Tree of Thought (ToT) Yao et al. (2024) bolster these models’ reasoning capabilities, potentially reducing hallucinations. Integrating knowledge graphs enhances understanding of factual information and concept relationships, aiding content generation and fact-checking. Specialized verification models cross-reference outputs against curated knowledge to identify inaccuracies, while bias detection and mitigation techniques promote fairness. Finally, ethical guidelines and regulatory frameworks governing the responsible use of curated knowledge in AI development mitigate risks and foster public trust, collectively improving the quality, accuracy, and trustworthiness of AI-generated content.
Multimodal Hallucination: Addressing hallucinations in multimodal large Foundation models requires a comprehensive approach spanning data-centric initiatives, cross-modal alignment efforts, architectural innovations, standardized benchmarking, reframing of hallucination, and enhancing interpretability and trust. Data-centric techniques for robust data collection, augmentation, and calibration ensure diverse and high-quality training data. Cross-modal alignment focuses on aligning representations across modalities through sophisticated architectures. Model architectural advancements involve designing specialized models capable of handling complex linguistic and visual inputs effectively. Establishing unified metrics and standardized benchmarks enables accurate assessment of hallucination and reliable performance evaluations. Reframing hallucination as a feature explores its integration into downstream applications, optimizing for the human experience. Finally, developing techniques for interpreting model behavior, visualizing internals, and improving reliability assessment fosters trust in MLLMs. This multifaceted approach collectively addresses the critical hallucination challenge, paving the way for more reliable and trustworthy multimodal AI systems.
9 Conclusion
This survey paper systematically categorizes existing research on hallucination within FMs, providing comprehensive insights into critical facets, including detection, mitigation, tasks, datasets, and evaluation metrics. It addresses the pervasive impact of hallucination in FMs, acknowledging its impact across diverse domains. By examining recent advancements in detection and mitigation techniques, the paper underscores the importance of addressing this challenge, given FMs’ indispensable role in critical tasks. Its primary contribution lies in introducing a structured taxonomy for classifying hallucination in FMs, spanning text, image, video, and audio domains.

### v4

@@
-Unveiling Hallucination in Text, Image, Video, and Audio Foundation Models: A Comprehensive Review
+A Comprehensive Survey of Hallucination in Large Language, Image, Video and Audio Foundation Models
 Abstract
@@
 By synthesizing recent advancements in detecting and mitigating hallucination across various modalities, the paper aims to provide valuable insights for researchers, developers, and practitioners.
-Essentially, it establishes a clear framework encompassing definition, taxonomy, and detection strategies for addressing hallucination in multimodal foundation models, laying the foundation for future research in this pivotal area.
+Essentially, it establishes a clear framework encompassing definition, taxonomy, and detection strategies for addressing hallucination in multimodal foundation models, laying the foundation for future research and development in this pivotal area.
 1 Introduction
@@
 1.1 Hallucination
-Hallucination refers to instances where FMs generate content that appears plausible and emulates human-like patterns but lacks a coherent understanding of the underlying context or factual grounding Xu et al.
+Hallucination refers to instances where FMs generate plausible but factually incoherent or absurd content lacking proper context understanding Xu et al.
 (2024b).
-These hallucinated outputs can manifest in various forms, ranging from trivial factual inaccuracies to more severe cases where the model generates entirely imaginary or absurd content.
-This issue is not limited to textual modalities but extends across diverse domains, including images, videos, and audio generated by large-scale foundation models.
-The root causes of hallucination are multifaceted, potentially stemming from biases in training data, limited access to current information, or the model’s inherent constraints in comprehending and generating contextually precise responses.
-Deploying these powerful models without addressing their hallucination tendencies can have severe consequences, perpetuating misinformation, leading to incorrect conclusions, and potentially causing adverse effects in critical applications.
-Addressing hallucinations in FMs has become an active research area and a key focus for development efforts, with ongoing efforts to mitigate this behavior.
-Strategies employed include fine-tuning models on domain-specific data, leveraging diverse training data to enhance model robustness, and developing improved evaluation metrics to identify and reduce hallucination tendencies.
+These hallucinated outputs can range from minor inaccuracies to completely imaginary content, manifesting across text, images, videos, and audio generated by large models.
+Root causes may include biases in training data, limited access to up-to-date information, or inherent model constraints in contextual comprehension and response generation.
+Deploying such powerful models without addressing hallucinations risks perpetuating misinformation, incorrect conclusions, and adverse effects in critical applications.
+Mitigating hallucinations has become an active research focus, with strategies like fine-tuning with domain-specific data, using diverse, robust training data, and developing improved evaluation metrics to identify and reduce hallucination tendencies.
 1.2 Types of Hallucination
-Hallucinations in large foundation models can manifest in various forms, each posing unique challenges and requiring tailored strategies for mitigation.
-These types of hallucinations can be categorized as follows: Contextual disconnection Zhang et al.
-(2023d) describes a situation in which the output or content produced by a model across different modalities is inconsistent or out of sync with the context that the user or the input data have provided or expected.
-Semantic distortion  Tjio et al.
+Hallucinations in large FMs can manifest in various forms and can be categorized as follows: Contextual disconnection: Zhang et al.
+(2023d) described a situation in which the output or content produced by a model across different modalities is inconsistent or out of sync with the context that the user or the input data provided or expected.
+Semantic distortion:  Tjio et al.
 (2022) refers to a type of inconsistency or error in generated content where the semantics or underlying meaning of the input is misrepresented or altered in the output.
-Content hallucination Moernaut et al.
-(2018)is the term used to describe a phenomenon seen in generative models when features or elements that are generated as output are either unreal given the context or absent from the input data.
-Factual inaccuracy  Zhang et al.
-(2023d) describes a kind of error seen in generative models when information that is inaccurate, deceptive, or at odds with the known facts appears in the generated output.
-Figure 1 illustrates various types of hallucinations.
-Figure 1: Types of hallucinations illustrated: This image shows several types of hallucinations.
-Proper explanations of hallucinations are indicated as hallucinated elements (HE) and are highlighted in bold red text.
+Content hallucination is the term used to describe a phenomenon seen in generative models when features or elements that are generated as output are either unreal given the context or absent from the input data Moernaut et al.
+(2018).
+Factual inaccuracy:  Zhang et al.
+(2023d) described a kind of error seen in generative models when information that is inaccurate, deceptive, or at odds with the known facts appears in the generated output.
+Figure 1 illustrates various types of hallucinations with examples.
+Figure 1: Illustration of Hallucination types. Proper explanations of hallucinations are indicated as hallucinated elements (HE) and are highlighted in bold red text.
 1.3 Motivation and Contributions
 Most of the existing survey papers have explored hallucination in the context of large language models (LLMs) Huang et al.
-(2023)Tonmoy et al.
+(2023)Chen et al.
+(2024b)Tonmoy et al.
 (2024).
@@
 (2024a)Sahoo et al.
-(2024)Rawte et al.
-(2023b).
-To address this gap, our survey aims to provide a holistic and multimodal perspective on the hallucination challenge in FMs.
-This review comprehensively examines existing research across language, vision, video, and audio domains to understand the mechanisms, detection methods, and mitigation strategies for hallucination in FMs.
+(2024b)Rawte et al.
+(2023).
+To address this gap, the present survey aims to provide a holistic and multimodal perspective on the hallucination challenge in FMs across language, vision, video, and audio domains.
 It serves as a vital resource for researchers and practitioners, aiding in the development of more robust AI solutions.
-Additionally, it includes a detailed taxonomy diagram in Figure.
-2 and a summarized Table 1 illustrating recent advancements across different modalities.
+Additionally, it includes a detailed taxonomy diagram in Fig.
+2 and a summarized Table 1 (refer to section 9.1 of the appendix) illustrating recent advancements across different modalities.
 The contributions of this survey paper are as follows:
+  • This survey provides a comprehensive overview of hallucination detection and mitigation techniques designed specifically for multimodal foundation models, extending beyond the typical focus solely on language models.
   • Establish a precise definition and structured taxonomy of hallucination in the context of large-scale foundation models.
-  • Identify the key factors and mechanisms that contribute to the emergence of hallucination across different modalities.
   • We have presented the various detection and mitigation strategies that have been proposed to address the hallucination problem in a multimodal setting.
-  • We have provided a comprehensive summary of the methodologies pertaining to hallucination techniques in large foundational models in Table 1, detailing their approaches to hallucination detection, mitigation, task considerations, datasets utilized, and evaluation metrics employed.
-This will offer readers a concise overview of recent advancements in this field.
-The paper is structured as follows: Section 2 delves into hallucinations in LLMs, while Section 3 explores large vision language models.
-Additionally, Section 4 and Section 5 examine hallucinations in large video and large audio models, respectively.
-Section 6 presents a discourse on the ramifications of hallucinations, exploring whether they pose advantages or disadvantages.
-Subsequently, Section 7 addresses the limitations of the current work, while Section 8 outlines potential future directions.
-Finally, Section 9 presents the concluding remarks.
+  • Highlight the open challenges and future research directions in this critical area.
 {forest}
-forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=center, text centered, minimum width=5em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style= rotate=90, child anchor=north, parent anchor=south, anchor=center , , where level=1text width=15em,font=,, where level=2text width=14em,font=,, where level=3minimum width=10em,font=,, where level=4text width=26em,font=,, where level=5text width=20em,font=,, [ Hallucinations, for tree=fill=bg16, text width=14em [ Text §2, for tree=fill=red!50, text width=18em [ Detection, for tree=fill=bg1, text width=13.5em [ FACTOID Rawte et al.
-(2024b), for tree=fill=pg56,text width=28em ] [ FACTOR Muhlgay et al.
+forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=center, text centered, minimum width=5em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=15em,font=,, where level=2text width=14em,font=,, where level=3minimum width=10em,font=,, where level=4text width=26em,font=,, where level=5text width=20em,font=,, [ Hallucinations, for tree=fill=paired-light-red!80, text width=14em [ Text §2, for tree=fill=red!50, text width=18em [ Detection
+[ FACTOR Muhlgay et al.
 (2023), for tree=fill=pg56,text width=28em ] [ FactCHD Chen et al.
-(2023b), for tree=fill=pg56,text width=28em ] [ HalluQA Cheng et al.
+(2023b), for tree=fill=pg56,text width=28em ] ] [ Inference Classifier [ HalluQA Cheng et al.
 (2023), for tree=fill=pg56,text width=28em ] [ HaluEval Li et al.
-(2023b), for tree=fill=pg56,text width=28em ] [ SelfCheckGPT Manakul et al.
-(2023), for tree=fill=pg56,text width=28em ] [ FACTSCORE Min et al.
+(2023b), for tree=fill=pg56,text width=28em ] ] [ Self-Evaluation [ SelfCheckGPT Manakul et al.
+(2023), for tree=fill=pg56,text width=28em ] ] [ Evidence Retrieval [ FACTSCORE Min et al.
 (2023), for tree=fill=pg56,text width=28em ] [ FacTool Chern et al.
-(2023), for tree=fill=pg56,text width=28em ] ] [ Mitigation, for tree=fill=bg2, text width=13.5em [ DoLa Chuang et al.
-(2023), for tree=fill=bg4,text width=28em ] [ CoK Li et al.
-(2023d), for tree=fill=bg4,text width=28em ] [ MixAlign Knowledge , for tree=fill=bg4,text width=28em ] [ SELF-FAMILIARITY Luo et al.
-(2023), for tree=fill=bg4,text width=28em ] [ CoVe Dhuliawala et al.
-(2023), for tree=fill=bg4,text width=28em ] [ HALOCHECK Elaraby et al.
-(2023), for tree=fill=bg4,text width=28em ] [ Instructional Prompting Varshney et al.
-(2023), for tree=fill=bg4,text width=28em ] [ PURR Chen et al.
-(2023a), for tree=fill=bg4,text width=28em ] [ LLM-AUGMENTER Peng et al.
-(2023), for tree=fill=bg4,text width=28em ] ] ] [ Image §3, for tree=fill=orange!50, text width=18em [ Detection, for tree=fill=bg1, text width=13.5em [ HallusionBench Guan et al.
-(2023), for tree=fill=dandyshade,text width=28em ] [ GAVIE Liu et al.
-(2023), for tree=fill=dandyshade,text width=28em ] [ PhD Liu et al.
-(2024b), for tree=fill=dandyshade,text width=28em ] [ VHTest Huang et al.
-(2024), for tree=fill=dandyshade,text width=28em ] [ M-HalDetect Gunjal et al.
-(2024), for tree=fill=dandyshade,text width=28em ] [ ChartBench Xu et al.
-(2023b), for tree=fill=dandyshade,text width=28em ] [ MSG-MCQ Lu et al.
-(2024), for tree=fill=dandyshade,text width=28em ] [ MMVP Tong et al.
-(2024), for tree=fill=dandyshade,text width=28em ] [ REVO-LION Liao et al.
-(2023), for tree=fill=dandyshade,text width=28em ] [ CIEM Hu et al.
-(2023), for tree=fill=dandyshade,text width=28em ] [ FAITHSCORE Jing et al.
-(2023), for tree=fill=dandyshade,text width=28em ] [ POPE Li et al.
-(2023f), for tree=fill=dandyshade,text width=28em ] [ VQA Changpinyo et al.
-(2022), for tree=fill=dandyshade,text width=28em ] [ HaELM Wang et al.
-(2023), for tree=fill=dandyshade,text width=28em ] [ NOPE Lovenia et al.
-(2023), for tree=fill=dandyshade,text width=28em ] [ MMHAL-BENCH Sun et al.
-(2023), for tree=fill=dandyshade,text width=28em ] [ TouchStone Bai et al.
-(2023), for tree=fill=dandyshade,text width=28em ] ] [ Mitigation, for tree=fill=bg2, text width=13.5em [ HalluciDoctor Yu et al.
-(2023a), for tree=fill=bg8,text width=28em ] [ LURE Zhou et al.
-(2023), for tree=fill=bg8,text width=28em ] [ MARINE Zhao et al.
-(2024), for tree=fill=bg8,text width=28em ] [ FDPO Gunjal et al.
-(2024), for tree=fill=bg8,text width=28em ] [ ViGoR Yan et al.
-(2024), for tree=fill=bg8,text width=28em ] [ HA-DPO Zhao et al.
-(2023), for tree=fill=bg8,text width=28em ] [ VIGC Wang et al.
-(2024), for tree=fill=bg8,text width=28em ] [ Data Centric Approach Lu et al.
-(2024), for tree=fill=bg8,text width=28em ] [ MoF Tong et al.
-(2024), for tree=fill=bg8,text width=28em ] [ InternLM-XComposer Zhang et al.
-(2023b), for tree=fill=bg8,text width=28em ] [ VCD Leng et al.
-(2023), for tree=fill=bg8,text width=28em ] [ Factually Augmented RLHF Sun et al.
-(2023), for tree=fill=bg8,text width=28em ] [ ObjMLM Dai et al.
-(2022), for tree=fill=bg8,text width=28em ] ] ] [ Video §4 , for tree=fill=bg5, text width=18em [ Detection, for tree=fill=bg1, text width=13.5em [ EMScore Shi et al.
-(2022), for tree=fill=bg7,text width=28em ] ] [ Mitigation, for tree=fill=bg2, text width=13.5em [ FactVC Liu and Wan (2023), for tree=fill=bg13,text width=28em ] [ CLearViD Chuang and Fazli (2023), for tree=fill=bg13,text width=28em ] [ MGAT He et al.
-(2022), for tree=fill=bg13,text width=28em ] ] ] [ Audio §5 , for tree=fill=lime!50, text width=18em [ Detection, for tree=fill=bg1, text width=13.5em [ PAM Deshmukh et al.
-(2024), for tree=fill=bg15,text width=28em ] [ CompA Ghosh et al.
-(2023), for tree=fill=bg15,text width=28em ] ] [ Mitigation, for tree=fill=bg2, text width=13.5em [ MusicLDM Chen et al.
-(2024), for tree=fill=bg14,text width=28em ] [ Re-AudioLDM Yuan et al.
-(2024), for tree=fill=bg14,text width=28em ] [ SECap Xu et al.
-(2024a), for tree=fill=bg14,text width=28em ] [ RECAP Ghosh et al.
-(2024), for tree=fill=bg14,text width=28em ] [ Cacophony Zhu and Duan (2024), for tree=fill=bg14,text width=28em ] [ EnCLAP Kim et al.
-(2024), for tree=fill=bg14,text width=28em ] ] ] ]
+(2023), for tree=fill=pg56,text width=28em ] ] ] [ Mitigation
+[ LLM-AUGMENTER Peng et al.
+(2023), for tree=fill=pg58,text width=28em ] ] [ Chain-of-Thought and Reasoning Enhancement [ CoK Li et al.
+(2023e), for tree=fill=pg58,text width=28em ] [ CoVe Dhuliawala et al.
+(2023), for tree=fill=pg58,text width=28em ] ] [ Alignment and Fine-Tuning [ MixAlign Knowledge , for tree=fill=pg58,text width=28em ] [ PURR Chen et al.
+(2023a), for tree=fill=pg58,text width=28em ] ] [ Self-Verification and Consistency Checking [ SELF-FAMILIARITY Luo et al.
+(2023), for tree=fill=pg58,text width=28em ] [ HALOCHECK Elaraby et al.
+(2023), for tree=fill=pg58,text width=28em ] ] [ Prompt Engineering and Instructional Techniques [ Instructional Prompting Varshney et al.
+(2023), for tree=fill=pg58,text width=28em ] ] ] ] [ Image §3, for tree=fill=orange!50, text width=18em [ Detection
+[ PhD Liu et al. (2024b), for tree=fill=Dandelion,text width=28em ] [ REVO-LION Liao et al. (2023), for tree=fill=Dandelion,text width=28em ] ] [ Visual Question Answering
+[ VQA Changpinyo et al. (2022), for tree=fill=Dandelion,text width=28em ] ] [ Benchmark and Evaluation Datasets
+[ ChartBench Xu et al.
+(2023b), for tree=fill=Dandelion,text width=28em ] [ MMHAL-BENCH Sun et al.
+(2023), for tree=fill=Dandelion,text width=28em ] [ TouchStone Bai et al.
+(2023), for tree=fill=Dandelion,text width=28em ] ] [ Multimodal Evaluation
+[ M-HalDetect Gunjal et al. (2024), for tree=fill=Dandelion,text width=28em ] ] [ Quantitative Metrics and Scoring Systems
+[ POPE Li et al.
+(2023g), for tree=fill=Dandelion,text width=28em ] [ NOPE Lovenia et al.
+(2023), for tree=fill=Dandelion,text width=28em ] ] [ CIEM Hu et al.
+(2023), for tree=fill=Dandelion,text width=28em ] [ MSG-MCQ Lu et al.
+(2024), for tree=fill=Dandelion,text width=28em ] [ HaELM Wang et al.
+(2023), for tree=fill=Dandelion,text width=28em ] ] [ Mitigation
+[ LURE Zhou et al. (2023), for tree=fill=lightgray,text width=28em ] [ VCD Leng et al. (2023), for tree=fill=lightgray,text width=28em ] ] [ Fine-Tuning and Post-Processing
+[ HA-DPO Zhao et al. (2023), for tree=fill=lightgray,text width=28em ] [ InternLM-XComposer Zhang et al. (2023b), for tree=fill=lightgray,text width=28em ] ] [ Generative Adversarial Networks
+[ MARINE Zhao et al. (2024), for tree=fill=lightgray,text width=28em ] ] [ Data-Centric Approaches
+[ Data Centric Approach Lu et al. (2024), for tree=fill=lightgray,text width=28em ] ] [ Multi-Modal Fusion and Object-Level Modeling
+[ ObjMLM Dai et al. (2022), for tree=fill=lightgray,text width=28em ] ] [ Reinforcement Learning
+] ] ] [ Video §4 , for tree=fill=bg25, text width=18em [ Detection
+] [ Mitigation
+[ CLearViD Chuang and Fazli (2023), for tree=fill=pg60,text width=28em ] [ MGAT He et al.
+(2022), for tree=fill=pg60,text width=28em ] ] ] [ Audio §5 , for tree=fill=lime!50, text width=18em [ Detection
+[ CompA Ghosh et al. (2023), for tree=fill=pg70,text width=28em ] ] [ Mitigation
+[ Re-AudioLDM Yuan et al. (2024), for tree=fill=red!50,text width=28em ] ] [ Retrieval-Based Approaches
+] [ Self-Supervised and Contrastive Learning
+[ EnCLAP Kim et al. (2024), for tree=fill=red!50,text width=28em ] ] [ Cacophony Zhu and Duan (2024), for tree=fill=red!50,text width=28em ] ] ] ]
 Figure 2: Taxonomy of hallucination in large foundation models, organized around detection and mitigation techniques.
@@
 Despite the progress of LLMs, a notable challenge persists in their proneness to hallucinate, impeding their practical implementation.
-For instance, the illustration in Figure 3 exemplifies the generated response by the LLM, showcasing indications of hallucination.
+For instance, the illustration in Figure 3 exemplifies the generated response by the LLM, showing indications of hallucination.
 2.1 Hallucination Detection and Mitigation
-Identifying hallucinations in LLMs is crucial for ensuring the credibility and reliability of their results, especially in scenarios requiring factual correctness.
+Detecting hallucinations in LLMs is crucial for ensuring the credibility and reliability of their results, especially in scenarios requiring factual correctness.
 Existing fact-checking methods often rely on complex modules or external databases, requiring either output probability distributions or interfacing with external sources.
-SelfCheckGPT Manakul et al.
+The SelfCheckGPT method Manakul et al.
 (2023) offers a zero-resource black-box solution for detecting hallucinations in any LLM without relying on external resources.
@@
 Continuing the exploration of methods for passage-level hallucination detection, Yang et al.
-(2023) proposed a novel self-check approach based on reverse validation, aiming to identify factual errors without external resources automatically.
+(2023) proposed a novel self-check approach based on reverse validation, aiming to automatically identify factual errors without external resources.
 They introduced a benchmark, Passage-level Hallucination Detection (PHD), generated using ChatGPT and annotated by human experts to assess different methods.
@@
 To address this, Min et al.
-(2023) introduced FACTSCORE (Factual Precision in Atomicity Score), a novel evaluation method that breaks down text into individual facts and measures their reliability.
+(2023) introduced FACTSCORE (Factual Precision in Atomicity Score), a new evaluation method that breaks down text into individual facts and measures their reliability.
 Huang and Chang (2023) introduced a unique strategy to mitigate hallucination risks in LLMs by drawing parallels with established web systems.
-They identified the absence of a "citation" mechanism, which refers to acknowledging or referencing sources or evidence, as a significant gap.
+They identified the absence of a "citation" mechanism in LLMs, which refers to acknowledging or referencing sources or evidence, as a significant gap.
 Addressing the need to identify factual inaccuracies in LLM-generated content, Rawte et al.
-(2024b) develop a multi-task learning (MTL) framework, integrating advanced long text embeddings like e5-mistral-7b-instruct, along with models such as GPT-3, SpanBERT, and RoFormer.
-This MTL approach exhibited a substantial performance gain, achieving an average accuracy improvement of 40% on the FACTOID benchmark when compared to leading textual entailment methods.
+(2024b) developed a multi-task learning (MTL) framework, integrating advanced long text embeddings like e5-mistral-7b-instruct, along with models such as GPT-3, SpanBERT, and RoFormer.
+This MTL approach demonstrated a 40% average improvement in accuracy on the FACTOID benchmark compared to leading textual entailment methods.
 Hallucination mitigation efforts have predominantly relied on empirical methods, leaving uncertainty regarding the possibility of complete elimination.
 To tackle this challenge, Xu et al.
-(2024b) introduces a formal framework defining hallucination as discrepancies observed between computable LLMs and a ground truth function.
-Through this framework, the study examines existing hallucination mitigation strategies and their practical implications for real-world LLM deployment.
+(2024b) introduced a formal framework defining hallucination as inconsistencies between computable LLMs and a ground truth function.
+The study examines existing hallucination mitigation strategies and their practical implications for real-world LLM deployment through this framework.
 Rawte et al.
 (2024c) introduced the Sorry, Come Again (SCA) prompting technique to address hallucination in contemporary LLMs.
-SCA enhances comprehension by employing optimal paraphrasing and injecting [PAUSE] tokens to delay LLM generation.
-It analyzes linguistic nuances in prompts and their impact on the hallucinated generation, highlighting the difficulties posed by prompts characterized by lower readability, formality, or concreteness.
-Rawte et al.
-(2023a) investigate how LLMs respond to factually correct and incorrect prompts, categorizing their hallucinations into mild, moderate, and alarming subcategories.
-Additionally, the paper introduced the Hallucination eLiciTation dataset, comprising 75,000 text snippets annotated by humans, and introduced a novel Hallucination Vulnerability Index metric.
-2.2 Domain Specific Works
-Hallucinations pose severe risks in critical fields such as healthcare, finance, and law.
-Reliability and accuracy are crucial in these sectors, as any form of hallucination can lead to significant and adverse consequences.
-2.2.1 Healthcare
-In response to the hallucinations in the medical domain LLMs, Pal et al.
-(2023) introduced the Medical Domain Hallucination Test (Med-HALT), a specialized benchmark dataset aimed at assessing and mitigating hallucinations.
-Med-HALT consists of a varied multinational dataset drawn from medical records spanning numerous countries, encompassing a total of seven datasets.
-Ahmad et al.
-(2023) outlined essential steps for creating dependable, trustworthy, and unbiased models, emphasizing the need for quantifying, validating, and mitigating hallucinations within the healthcare context.
-Ji et al.
-(2023) introduced an interactive self-reflection approach aimed at improving the accuracy and coherence of answers generated by medical question-answering systems using LLMs.
-Through knowledge acquisition and feedback on answer generation, this methodology enhances the factuality, consistency, and logical progression of responses.
-2.2.2 Finance
-An empirical investigation explored the inclination of LLMs to generate hallucinations while engaging in financial activities.
-Kang and Liu (2023) conducted an empirical investigation into the hallucination tendencies of LLMs in financial tasks.
-Their study assessed LLMs’ proficiency in explaining financial concepts, querying historical stock prices, and examined the efficacy of methods like few-shot learning and prompt-based tool learning in mitigating hallucination.
-Roychowdhury et al.
-(2023) proposed a novel Langchain-based approach aimed at transforming data tables into hierarchical textual data chunks, facilitating versatile financial question answering.
-The framework involves classifying user queries by intention, retrieving relevant data chunks, generating customized LLM prompts, and evaluating responses for hallucinations and confidence.
-2.2.3 Legal
-The conventional approach to abstractive text summarization typically employs an encoder-decoder architecture, wherein the encoder encapsulates the essence of the source text while the decoder generates the summary.
-However, this method may produce summaries containing irrelevant or inaccurate information, which poses a significant concern in legal contexts where accuracy is crucial.
-To address these issues, Feijo and Moreira (2023) introduced LegalSumm, which creates distinct "views" of the source text, trains summarization models to produce independent summaries, and employs an entailment module to assess their fidelity to the source.
-Deroy et al.
-(2023) investigated the readiness of LLMs for generating abstractive summaries of case judgments by applying SOTA models to Indian court cases.
-While abstractive models generally scored slightly higher, the authors noted inconsistencies and hallucinations in the generated summaries.
-Understanding the meaning of open-ended legal terms is very important for legal professionals.
-They often look at how such terms were used and interpreted in previous court cases.
-Savelka et al.
-(2023) evaluate GPT-4’s performance in generating factually accurate, clear, and relevant explanations of legal terms in legislation.
-A comparison is made between a baseline approach, where GPT-4 directly explains a legal term, and an augmented approach employing a legal information retrieval module to provide contextual sentences from case law.
-Dahl et al.
-(2024) presents the initial evidence on the frequency and types of inaccuracies in the legal domain, providing valuable insights for evaluating LLMs in legal contexts.
-By examining the structured format of American case law, the study assesses three major LLMs: GPT-3.5, PaLM 2, and Llama.
-2.3 Benchmark Evaluation
-In certain instances, LLMs engage in a phenomenon termed "hallucination snowballing," where they fabricate false claims to rationalize prior hallucinations despite acknowledging their inaccuracy.
+SCA enhances comprehension through optimal paraphrasing and injecting [PAUSE] tokens to delay LLM generation.
+It analyzes linguistic nuances in prompts and their impact on the hallucinated generation, emphasizing how prompts with lower readability, formality, or concreteness pose challenges.
+Benchmark Evaluation: In certain instances, LLMs engage in a phenomenon termed "hallucination snowballing," where they fabricate false claims to rationalize prior hallucinations despite acknowledging their inaccuracy Li et al.
+(2023c).
 To empirically explore this phenomenon, Zhang et al.
-(2023a) devise three question-answering datasets spanning diverse domains, wherein ChatGPT and GPT-4 often furnish inaccurate answers alongside explanations featuring at least one false claim.
-Significantly, the study suggested that the language model can discern these false claims as incorrect.
+(2023a) devised three question-answering datasets spanning diverse domains, wherein ChatGPT and GPT-4 often furnish inaccurate answers alongside explanations featuring at least one false claim.
+Significantly, the study suggests that the language model can discern these false claims as incorrect.
 Another benchmark dataset, FactCHD Chen et al.
 (2023b), was introduced to detect fact-conflicting hallucinations within intricate inferential contexts.
-It encompasses a range of datasets capturing different factual patterns and integrates fact-based evidence chains to improve assessment accuracy.
-Li et al.
-(2023b) introduced a dataset to assess the capability of LLMs to identify and recognize hallucinated or incorrect information.
-The outcomes highlighted ChatGPT’s inclination to produce hallucinated content, particularly on certain topics, introducing unverifiable information.
-Figure 4: Example of types of Vision Hallucination.
-Liu et al.
-(2024b) including visuals and the matching question-answer pairs and hallucination elements.
-While red annotated words do not exist or do not correspond to elements within the image, whereas words highlighted in green have correspondences within the image.
-Question and Answer are denoted by the letters Q and A, respectively.
+It encompasses a range of datasets capturing different factuality patterns and integrates fact-based evidence chains to improve assessment accuracy.
+Figure 4: Four IVL-Hallu examples in Prompted Hallucination Dataset(PhD)  Liu et al.
+(2024b) including visuals and the matching question-answer pairs and hallucination elements (HE).
+While words annotated in red do not exist or do not match within the image, words annotated in green have correspondences within the image.
+Question, Answer, and Statement are denoted by the letters Q, A, and S, respectively.
 3 Hallucination in Large Vision-Language Models
-Large Vision-Language Models (LVLMs) have garnered significant attention in the AI community for their capacity to handle visual and textual data simultaneously.
+Large Vision-Language Models (LVLMs) have garnered significant attention in the AI community for their ability to handle visual and textual data simultaneously Ghosh et al.
+(2024c)Ghosh et al.
+(2024b)Sahoo et al.
+(2024a)Ghosh et al.
+(2024d)Ghosh et al.
+(2024a).
 Nonetheless, similar to LLMs, LVLMs also confront the issue of hallucination.
@@
 Dai et al.
-(2022) investigate the issue of object hallucinations in Vision-Language Pre-training (VLP) models, where textual descriptions generated by these models contain non-existent or inaccurate objects based on input images.
+(2022) investigated the issue of object hallucinations in Vision-Language Pre-training (VLP) models, where textual descriptions generated by these models contain non-existent or inaccurate objects based on input images.
 Li et al.
-(2023f) reveal widespread and severe object hallucination issues and suggest that visual instructions may influence hallucination.
-They observe that objects frequently depicted in visual instructions or co-occurring with image objects are more susceptible to hallucination.
-To enhance the evaluation process of object hallucination, the authors introduced a polling-based query method called POPE, which demonstrates improved stability and flexibility in assessing object hallucination.
+(2023g) revealed widespread and severe object hallucination issues and suggested that visual instructions may influence hallucination, noting that objects frequently appearing in visual instructions or co-occurring with image objects are more likely to be hallucinated.
+To enhance the evaluation of object hallucination, they introduced a polling-based query method called POPE, which demonstrates improved stability and flexibility in assessing object hallucination.
 The absence of a standardized metric for assessing object hallucination has hindered progress in understanding and addressing this issue.
 To address this gap, Lovenia et al.
-(2023) introduced NOPE (Negative Object Presence Evaluation), a benchmark for evaluating object hallucination in vision-language models (VLMs) through visual question answering (VQA).
-Utilizing LLMs, the study generated a dataset of 29.5k synthetic negative pronoun (NegP) instances for NOPE.
-It thoroughly assessed the capability of 10 VLMs in detecting the absence of objects in visual questions, in addition to their typical performance on visual questions across nine other VQA datasets.
-Existing research focuses primarily on object hallucination, overlooking the LVLM hallucinations.
+(2023) introduced NOPE (Negative Object Presence Evaluation), a novel benchmark for evaluating object hallucination in vision-language (VL) models through visual question answering (VQA).
+Utilizing LLMs, the study generates 29.5k synthetic negative pronoun (NegP) data for NOPE.
+It extensively evaluates the performance of 10 VL models in discerning the absence of objects in visual questions, alongside their standard performance on visual questions across nine other VQA datasets.
+Most existing efforts focused primarily on object hallucination, overlooking the diverse types of LVLM hallucinations.
 Liu et al.
-(2024b) delve into Intrinsic Vision-Language Hallucination (IVL-Hallu) and proposes several novel IVL-Hallu tasks, including attribute, object, multi-modal conflicting, and counter-common-sense hallucination.
-They introduced a challenging benchmark dataset to assess and explore IVL-Hallu, conducting experiments on five LVLMs that revealed their limited effectiveness in addressing the proposed tasks.
+(2024b) delved into Intrinsic Vision-Language Hallucination (IVL-Hallu) and proposed several novel IVL-Hallu tasks categorized into four types: attribute, object, multi-modal conflicting, and counter-common-sense hallucination.
+To assess and explore IVL-Hallu, they introduced a challenging benchmark dataset and conducted experiments on five LVLMs, revealing their incapacity to effectively address the proposed IVL-Hallu tasks.
 To mitigate object hallucination in LVLMs without resorting to costly training or API reliance, Zhao et al.
-(2024) introduced MARINE, a training-free and API-free solution.
-MARINE enhances LVLMs’ visual comprehension by combining existing open-source vision models and leveraging guidance without classifiers to incorporate object grounding features, thereby enhancing the precision of generated outputs.
-Evaluations across six LVLMs revealed MARINE’s effectiveness in reducing hallucinations and enhancing output detail, validated through assessments using GPT-4V.
+(2024) introduced MARINE, which is both training-free and API-free.
+MARINE enhances the visual understanding of LVLMs by integrating existing open-source vision models and utilizing guidance without classifiers to integrate object grounding features, thereby improving the precision of the generated outputs.
+Evaluations across six LVLMs reveal MARINE’s effectiveness in reducing hallucinations and enhancing output detail, validated through assessments using GPT-4V.
+Deng et al.
+(2024) introduced a CLIP-Guided Decoding (CGD) training-free approach to reduce object hallucination at decoding time.
 HalluciDoctor Yu et al.
 (2023a) tackled hallucinations in Multi-modal Large Language Models (MLLMs) by using human error detection to identify and eliminate various types of hallucinations.
-Through rebalancing data distribution via counterfactual visual instruction expansion, they successfully mitigate 44.6% of hallucinations while maintaining competitive performance.
+By rebalancing data distribution via counterfactual visual instruction expansion, they successfully mitigate 44.6% of hallucinations while maintaining competitive performance.
 Despite proficiency in visual semantic comprehension and meme humor, MLLMs struggle with chart analysis and understanding.
 Addressing this, Xu et al.
-(2023b) propose ChartBench, a benchmark assessing chart comprehension.
+(2023b) proposed ChartBench, a benchmark assessing chart comprehension.
 ChartBench exposes MLLMs’ limited reasoning with complex charts, prompting the need for novel evaluation metrics like Acc+ and a handcrafted prompt, ChartCoT.
@@
 The performance of InternLM-XComposer’s text-image composition is evaluated through a robust procedure involving both human assessment and comparison to GPT4-Vision, with the model demonstrating competitive performance against solutions like GPT4-V and GPT3.5.
-Cutting-edge LVLMs like InstructBLIP Dai et al.
-(2024), while generating visually grounded responses, often include inaccuracies like fictional items and flawed relationships.
-To enhance accuracy, Gunjal et al.
-(2024) introduced M-HalDetect, a pioneering multi-modal fine-grained hallucination detection dataset.
-This dataset serves as a benchmark for training LVLMs, leading to more precise outputs.
-Using fine-grained multi-modal reward models and enhancing FDPO significantly reduced InstructBLIP’s hallucination rate.
-These methods not only improve accuracy across LVLMs like LLaVA and mPLUG-OWL but also highlight the versatility and effectiveness of M-HalDetect in identifying and minimizing hallucinations.
-Despite advancements in multi-modal tasks, LMMs often generate descriptions that are inconsistent with the accompanying image or human instructions.
-To address this, Liu et al.
-(2023) develop LRV-Instruction, a comprehensive dataset comprising 400k visual instructions across 16 tasks.
-This dataset includes both positive and negative instructions in various styles and semantic levels.
-Through LRV-Instruction, the issue of hallucination in existing LMMs was extensively examined, confirming its effectiveness in enhancing visual instruction tuning.
-Additionally, they introduced GAVIE, a novel method for evaluating visual instruction tuning without the need for human-labeled answers, which can be adapted to different types of instructions.
-The LVLM Hallucination Revisor (LURE) algorithm Zhou et al.
-(2023) was developed to correct object hallucination in LVLMs by refining the descriptions to produce more accurate and less hallucinatory outputs.
-Its methodology is based on an in-depth statistical analysis that identifies key factors contributing to object hallucination, such as the co-occurrence of certain objects in images, the uncertainty associated with objects during LVLM decoding, and the tendency for hallucinations to occur towards the end of the generated text.
-LURE is designed for seamless integration with a variety of LVLMs.
-When tested across multiple LVLMs, the integration of LURE resulted in significant enhancements in object hallucination correction, consistently outperforming other methods in both GPT and human evaluations based on various metrics.
+Wang et al.
+(2024b) proposed the Instruction Contrastive Decoding (ICD) method to reduce hallucinations during LVLM inference.
+A recent study Huang et al.
+(2024a) proposed a novel decoding approach that introduces an over-confidence penalty and a retrospective allocation strategy to mitigate hallucination issues without requiring additional data or retraining.
 3.2 Benchmark Evaluation
@@
 To address this, Li et al.
-(2023e) propose a novel data collection approach that synthesizes images and dialogues synchronously for visual instruction tuning, yielding a large dataset of image-dialogue pairs and multi-image instances.
+(2023f) proposed a novel data collection approach that synthesizes images and dialogues synchronously for visual instruction tuning, yielding a large dataset of image-dialogue pairs and multi-image instances.
 Huang et al.
-(2024) introduced VHTest, a benchmark dataset with 1,200 diverse visual hallucinations (VH) instances across 8 VH modes.
+(2024b) introduced VHTest, a benchmark dataset with 1,200 diverse visual hallucinations (VH) instances across 8 VH modes.
 Evaluation of three SOTA MLLMs showed varying performance, with GPT-4V exhibiting lower hallucination than MiniGPT-v2.
@@
 They proposed three main categories of methods to mitigate hallucination: data-driven approaches, training adjustments, and post-processing techniques.
-Additionally, Wang et al.
-(2024) proposed the Visual Instruction Generation and Correction (VIGC) framework to address the shortage of high-quality instruction-tuning data for MLLMs.
+Wang et al.
+(2024a) proposed the Visual Instruction Generation and Correction (VIGC) framework to address the scarcity of high-quality instruction-tuning data for MLLMs.
 VIGC enables MLLMs to generate diverse instruction-tuning data while iteratively refining its quality through Visual Instruction Correction (VIC), mitigating hallucination risks.
 The framework produces diverse, high-quality data for fine-tuning models, validated through evaluations, improving benchmark performance, and overcoming language-only data limitations.
-Figure 5: A video featuring descriptions generated by VLTinT model and ground truth (GT) with description errors highlighted in red italics Chuang and Fazli (2023).
+Figure 5: A video featuring descriptions generated by VLTinT model and ground truth (GT) with description errors highlighted in red italics. Chuang and Fazli (2023).
 4 Hallucinations in Large Video Models
-Large Video Models (LVMs) represent a significant advancement, allowing for the processing of video data at scale.
+Large video models (LVMs) represent a significant advancement, allowing for processing video data at scale.
 Despite their potential for various applications like video understanding and generation, LVMs face challenges with hallucinations, where misinterpretations of video frames can result in artificial or inaccurate visual data.
-This issue arises due to the video data complexity, which requires the model to process and comprehend it thoroughly.
 Figure 5 demonstrates the instances of hallucination observed in LVMs.
@@
 To address this, Mun et al.
-(2019) introduces a novel approach to modeling temporal dependencies and leveraging context for coherent storytelling.
+(2019) introduced a novel approach to modeling temporal dependencies and leveraging context for coherent storytelling.
 By integrating an event sequence generation network and a sequential video captioning network trained with reinforcement learning and two-level rewards, the model captures contextual information more effectively, yielding coherent and accurate captions while minimizing the risk of hallucinations.
-Liu and Wan (2023) introduces a novel weakly-supervised, model-based factuality metric called FactVC, which outperforms previous metrics.
-Furthermore, they provide two annotated datasets to promote further research in assessing the factuality of video captions.
+Liu and Wan (2023) introduced a novel weakly-supervised, model-based factuality metric called FactVC, which outperforms previous metrics.
+Furthermore, they provided two annotated datasets to promote further research in assessing the factuality of video captions.
 Wu and Gao (2023) proposed a context-aware model that incorporates information from past and future events to influence the description of the current event conditionally.
 Their approach utilizes a robust pre-trained context encoder to encode information about the surrounding context events, which is then integrated into the captioning module using a gate-attention mechanism.
-Experimental findings on the YouCookII and ActivityNet datasets demonstrate that the proposed context-aware model outperforms existing context-aware and pre-trained models significantly.
+Experimental findings on the YouCookII and ActivityNet datasets demonstrate that the proposed context-aware model outperforms existing context-aware and pre-trained models by a significant margin.
 To enhance dense video captioning,  Zhou et al.
-(2024) introduces a streaming model comprising a memory module for long video handling and a streaming decoding algorithm enabling predictions before video completion.
-This approach notably boosts performance on prominent dense video captioning benchmarks, such as YouCook2, ActivityNet, and ViTT.
+(2024) introduced a streaming model comprising a memory module for long video handling and a streaming decoding algorithm enabling predictions before video completion.
+This approach notably boosts performance on dense video captioning benchmarks such as ActivityNet, YouCook2, and ViTT.
 Video infilling and prediction tasks are crucial for assessing a model’s ability to comprehend and anticipate the temporal dynamics within video sequences Höppe et al.
@@
 Yu et al.
-(2023b) aims to tackle the opposite issue rather than depending on using pixels from other frames.
+(2023b) aimed to tackle the opposite issue rather than depending on using pixels from other frames.
 The suggested method presents a Deficiency-aware Masked Transformer (DMT), a dual-modality-compatible inpainting framework.
@@
 Kulal et al.
-(2023) introduces a method for realistically inserting people into scenes.
+(2023) introduced a method for realistically inserting people into scenes.
 The model seamlessly integrates individuals into scenes by deducing realistic poses based on the context and ensuring visually pleasing compositions.
-Chuang and Fazli (2023) introduces CLearViD, a transformer-based model that utilizes curriculum learning techniques to enhance performance.
+Chuang and Fazli (2023) introduced CLearViD, a transformer-based model that utilizes curriculum learning techniques to enhance performance.
 By adopting this approach, the model acquires more robust and generalizable features.
@@
 Zhang et al.
-(2006) created an innovative two-level hierarchical fusion method to hallucinate facial expression sequences from training video samples using only one frontal face image with a neutral expression.
+(2006) created a novel two-level hierarchical fusion method to hallucinate facial expression sequences from training video samples using only one frontal face image with a neutral expression.
 To effectively train the system, they introduced a dataset specifically designed for facial expression hallucination, which included 112 video sequences covering four types of facial expressions (happy, angry, surprise, and fear) from 28 individuals, resulting in the generation of reasonable facial expression sequences in both the temporal and spatial domains with less artifact.
@@
 Li et al.
-(2023c) introduced "VideoChat", a novel approach integrating video foundation models and LLMs through a learnable neural interface to enhance spatiotemporal reasoning, event localization, and causal relationship inference in video understanding.
+(2023d) introduced "VideoChat," a novel approach integrating video foundation models and LLMs through a learnable neural interface to enhance spatiotemporal reasoning, event localization, and causal relationship inference in video understanding.
 The researchers constructed a video-centric instruction dataset with detailed descriptions and conversations, emphasizing spatiotemporal reasoning and causal relationships.
@@
 (2023) curated a dataset of 2.4M video clips, showcasing a variety of plausible poses that align with the scene context.
-Figure 6: Audio hallucination examples for each classes - Type A: Involving hallucinations of both objects and actions $|$ Type B: Featuring accurate objects but hallucinated actions $|$ Type C: Displaying correct actions but hallucinated objects Nishimura et al.
+Figure 6: Audio hallucination examples for each classes - Type A: Involving hallucinations of both objects and actions: Type B: Featuring accurate objects but hallucinated actions; Type C: Displaying correct actions but hallucinated objects Nishimura et al.
 (2024).
@@
 (2023).
-Although these models have demonstrated remarkable capabilities across various domains, they are susceptible to hallucinations.
-These anomalies can take several forms, from creating unrealistic audio by piecing together fabricated snippets to injecting false information, such as quotes or facts, into summaries.
+While demonstrating remarkable capabilities, LAMs are susceptible to hallucinations – anomalies ranging from generating unrealistic audio by combining fabricated snippets to injecting false information like quotes or facts into summaries.
 Additionally, they may fail to accurately capture the inherent features of audio signals, such as timbre, pitch, or background noise Shen et al.
 (2023).
+Figure 6 presents one example of audio hallucinations.
 5.1 Hallucination Detection and Mitigation
 In the realm of audio captioning, where natural language descriptions for audio clips are automatically generated, a significant challenge arises from the over-reliance on the visual modality during the pre-training of audio-text models.
-This reliance introduced data noise and hallucinations, ultimately undermining the accuracy of the resulting captions.
+This reliance introduces data noise and hallucinations, ultimately undermining the accuracy of the resulting captions.
 To address this issue, Xu et al.
-(2023a) introduces an AudioSet tag-guided model designed to bootstrap large-scale audio-text data (BLAT).
+(2023a) introduced an AudioSet tag-guided model designed to bootstrap large-scale audio-text data (BLAT).
 Notably, this model sidesteps the incorporation of video, thus minimizing noise associated with the visual modality.
@@
 SECap Xu et al.
-(2024a), a framework designed for speech emotion captioning, aims to capture the intricate emotional nuances of speech using natural language.
+(2024a), a framework designed for speech emotion captioning.
+It aims to capture the intricate emotional nuances of speech using natural language.
 SECap utilizes various components, including LLaMA as the text decoder, HuBERT as the audio encoder, and Q-Former as the Bridge-Net, to generate coherent emotion captions based on speech features.
@@
 To address this, Elizalde et al.
-(2024) introduces the Contrastive Language-Audio Pretraining (CLAP) model.
+(2024) introduced the Contrastive Language-Audio Pretraining (CLAP) model.
 Pre-trained with 4.6 million diverse audio-text pairs, CLAP features a dual-encoder architecture, enhancing representation learning for improved task generalization across sound, music, and speech domains.
@@
 (2023) introduced LP-MusicCaps, a comprehensive dataset comprising 0.5 million audio clips accompanied by approximately 2.2 million captions.
-Leveraging LLMs, they train a transformer-based music captioning model with the dataset and assess its performance under zero-shot and transfer-learning scenarios, demonstrating its superiority over supervised baseline models.
+Leveraging LLMs, they trained a transformer-based music captioning model with the dataset and assessed its performance under zero-shot and transfer-learning scenarios, demonstrating its superiority over supervised baseline models.
 Nishimura et al.
@@
 6.
-In their investigation, they gathered 1000 sentences by soliciting audio information and then annotated them to determine whether they contained auditory hallucinations, further categorizing the type of hallucination if detected.
+They gathered 1000 sentences by soliciting audio information and then annotated them to determine whether they contained auditory hallucinations.
 To assess compositional reasoning in LAMs, Ghosh et al.
-(2023) introduced CompA, consisting of two expert-annotated benchmarks primarily focused on real-world audio samples.
-This benchmark is employed to fine-tune CompA-CLAP with a novel learning approach, enhancing its compositional reasoning skills and demonstrating substantial improvement over all the baseline models in tasks requiring compositional reasoning.
-TEXT Paper Detection Mitigation Task Dataset(s) Evaluation Metric(s)
-Manakul et al. (2023) Yes No QA Wikibio Entropy
-Li et al. (2022) Yes Yes QA, Dialog summarization Halueval Automatic
-Mündler et al. Yes Yes Text generation Manual F1 Score
-Chen et al. (2023a) No Yes Editing for attribution MCQ, Dialog Attribution, Preservation
-Zhang et al. (2023c) No Yes Question knowledge alignment Fuzzy QA Attributable to Identified Sources
-Zhang et al. (2023a) Yes No QA Manual Accuracy
-Peng et al. (2023) No Yes Task-oriented dialog News, Customer service F1 Score, Bleu-4
-Cui et al. (2023) No Yes QA Manual Ranking
-Azaria and Mitchell (2023) Yes No Classification Manual Accuracy
-Li et al. (2023d) Yes Yes Knowledge-intensive tasks Fever, QA Accuracy
-Elaraby et al. (2023) Yes Yes Consistency, Actuality, QA Manual NBA domain Pearson Correlation Coefficient
-Varshney et al. Yes Yes Text generation Wikibio Percentage of mitigated hallucination
-Jha et al. (2023) Yes No Dialog N/A N/A
-Pal et al. (2023) No No Reasoning hallucination Med-Halt Accuracy, Pointwise Score
-McKenna et al. (2023) Yes No Textual entailment Altered Directional Inference Entailment Probability
-Guerreiro et al. (2023) Yes Yes MT FLores 101, WMT ,TICO BLEU
-Huang and Chang (2023) Yes Yes N/A N/A N/A
-Luo et al. (2023) Yes Yes Concept extraction Concept-7 AUC, Accuracy, F1 Score
-Gao et al. (2022) Yes Yes Editing attribution NQ, SQA Auto-AIS (Attr_auto)
-Yang et al. (2023) Yes No
-Detect factual
-errors automatically
-PHD, WikiBio-GPT3
-Precision, Recall,
-F1 Score, Accuracy
-Min et al. (2023) Yes Yes Fact verification Manual(Wikipedia) FActScore
-Rawte et al. (2024b) Yes Yes Factual inaccuracies detection FACTOID HV I_auto
-Ahmad et al. (2023) Yes Yes Hallucination in healthcare N/A FActScores
-Ji et al. (2023) Yes Yes Generative and knowledge-intensive
-PubMedQA, MEDIQA2019,
-MedQuAD, and MASH-QA
-Unigram F1, ROUGE-L,
-Med-NLI, and CTRLEval
-Kang and Liu (2023) Yes Yes Hallucination in finance N/A FActScores
-Roychowdhury (2024) No Yes QA N/A N/A
-Savelka et al. (2023) No Yes Factual evaluation in legislation N/A N/A
-Dahl et al. (2024) Yes No Legal hallucination Manual N/A
-Rawte et al. (2024c) No Yes Comprehension enhancement SCA-90K Cosine similarity
-IMAGE Li et al. (2023f) Yes No Evaluation of object hallucination MSCOCO CHAIR, POPE
-Gunjal et al. (2024) Yes Yes VQA M-Hall Detect Accuracy
-Dai et al. (2022) No Yes Image captioning CHAIR CIDEr
-Lovenia et al. (2023) Yes No Object hallucination NOPE
-METEOR, Exact match accuracy,
-NegP Accuracy
-Liu et al. (2024b) Yes No Intrinsic vision-language hallucination PhD Accuracy
-Zhao et al. (2024) Yes Yes Non-existing object hallucination MSCOCO CHAIR, POPE, GPT-4V, recall
-Huang et al. (2024) Yes No Visual hallucination YNQ, OEQ Accuracy
-Rawte et al. (2024a) Yes No Video captioning ActivityNet-Fact, YouCook2-Fact FactVC
-Wang et al. (2024) No Yes
-Generate instruction data
-for vision-language
-VIGC-LLaVA-COCO,
-VIGC-LLaVA-Objects365
-Conv, Detail, Complex
-Yu et al. (2023a) Yes Yes
-Machine-generated
-visual instruction
-LLaVA-Instruction-158K CHAIR
-Guan et al. (2023) No Yes Visual questions HallusionBench Accuracy
-Liu et al. (2023) Yes Yes Vision language LRV-Instruction GAVIE
-Xu et al. (2023b) Yes No
-Evaluation of MLLMs on
-chart comprehension
-ChartBench Acc+
-Lu et al. (2024) Yes Yes Vision language MSG-MCQ Accuracy
-Tong et al. (2024) Yes No Visual question answering MMVP, VQA Accuracy
-Liao et al. (2023) Yes No Vision language REVO-LION Meta Quality (MQ)
-Hu et al. (2023) Yes Yes
-Visual captioning,
-Visual question answering
-CIEM
-Accuracy, Precision,
-Recall, F1 Score
-Jing et al. (2023) Yes No Meta-evaluation LLaVA-1k, MSCOCO-Cap FAITHSCORE
-Changpinyo et al. (2022) No Yes Multilingual visual question answering MaXM Accuracy
-Wang et al. (2023) Yes No Content generation N/A Precision, Recall, F1 Score
-Sun et al. (2023) No Yes Visual-language alignment MMHAL-BENCH N/A
-Bai et al. (2023) Yes No
-Evaluate hallucination of
-vision language model
-TouchStone Hallucination Score
-Zhou et al. (2023) No Yes Hallucination mitigation in LVMs MSCOCO CHAIR, BLEU, CLIP
-Yan et al. (2024) No Yes Visual grounding MMViG
-HL, CA, AA,
-RA, RL, RS, DL
-Zhao et al. (2023) Yes Yes Overcome hallucination in LVMs POPE, SHR Accuracy, Precision, F1 Score
-Zhang et al. (2023b) No Yes
-Image text comprehension
-and composition
-MMBench, SeedBench, QBench,
-MMBench-CN, Chinese Bench
-LR, AR, $R$,
-FP-C, FP-S, CP
-VIDEO Kulal et al. (2023) No Yes Affordance prediction Manual FID, FCKh
-Himakunthala et al. (2023) No Yes Video infilling, Scene prediction Manual N/A
-Li et al. (2023c) No Yes Visual dialogue Manual N/A
-Zhou et al. (2024) No Yes Video captioning
-ActivityNet Captions,
-YouCook2, ViTT
-CIDER, METEOR, SODAc
-Höppe et al. (2022) Yes No Video prediction BAIR, Kinetics 600, UCF-101 Frechet Video Distance
-Chuang and Fazli (2023) No Yes Video description Activity Net Captions, YouCook2
-METEOR, ROUGE_L, CIDER,
-BLEU_4, DIV-2, RE_4
-AUDIO Li et al. (2023a) No Yes Classification Manual Mean avg precision
-Doh et al. (2023) No Yes Audio captioning LP MusicCaps BLEU
-Xu et al. (2023a) No Yes Caption generation AudioCaps R@K, COCO & FENCE
-Liu and Wan (2023) No Yes Audio captioning MusciCaps BLEU
-Nishimura et al. (2024) Yes No Evaluation of LAMs LAION_CLAP,MS_CLAP Recall, Precision, F1 Score
-Table 1: Overview of the hallucination detection and mitigation landscape in FMs across modalities (Text, Image, Video, and Audio).
-Each work is categorized based on factors such as detection, mitigation, tasks, datasets, and evaluation metrics.
-6 Hallucination: Good or Bad?
-Hallucinations in large-scale models present a complex interplay between creativity and uncertainty.
-On one hand, the ability to traverse beyond conventional data boundaries can lead to the generation of novel and innovative outputs.
-Hallucinations can spark exploratory learning, revealing unexpected patterns and features within the data.
-They can also serve as a form of stress testing, improving the model’s robustness and adaptability.
-Furthermore, these unexpected outputs can even inspire human creativity, serving as a springboard for new ideas and perspectives Rawte et al.
-(2023b).
-However, this dual nature of hallucinations also introduced significant drawbacks.
-The quality and coherence of hallucinatory outputs can be questionable, posing challenges in applications where accuracy and reliability are paramount.
-Hallucinations can also propagate misinformation and biases present in the model’s training data, potentially reinforcing existing prejudices and eroding user trust.
-The reduced interpretability of these outputs can further undermine the model’s credibility and adoption.
-Ethical concerns arise when hallucinations produce inappropriate, offensive, or harmful content.
-Careful monitoring and control mechanisms are essential to prevent the generation of outputs that could cause harm or distress to users.
-Navigating this intricate balance between exploration and fidelity is crucial for maximizing the utility of large models while mitigating the risks associated with unexpected outputs.
-Overall, the phenomenon of hallucinations in large-scale models highlights the need for a nuanced understanding and strategic management of these capabilities.
-7 Limitations
+(2023) introduced CompA, consisting of two expert-annotated benchmarks and employed to fine-tune CompA-CLAP with a novel learning approach, leading to improved compositional reasoning abilities compared to baseline models on related tasks.
+6 Future Directions
+Researchers are actively investigating techniques to mitigate hallucinations, which is crucial for sensitive applications Tonmoy et al.
+(2024)Rawte et al.
+(2023).
+The Key future directions can include:
+Data Resources: Fine-tuning carefully curated high-quality data, integrating structured knowledge from knowledge graphs, and employing task/domain-specific alignment techniques to enhance accuracy and relevance.
+Automated Evaluation: Developing specialized metrics for factual accuracy and coherence, combining automated evaluation with human judgments, adversarial testing to identify weaknesses, and fine-tuning fact-checking datasets.
+Improving Detection and Mitigation: Leveraging reasoning mechanisms (e.g., Chain of Thought Wei et al.
+(2022), Tree of Thought Yao et al.
+(2024)), knowledge graph integration, specialized fact-checking models, bias mitigation techniques, active learning methodologies, and ethical guidelines/regulatory frameworks.
+Multimodal Hallucination: Data-centric initiatives, cross-modal alignment, architectural innovations, standardized benchmarking, reframing hallucination as a feature, and enhancing interpretability and trust for reliable multimodal AI systems.
+7 Conclusion
+This survey paper systematically categorizes existing research on hallucination within FMs, providing comprehensive insights into critical aspects such as detection, mitigation, tasks, datasets, and evaluation metrics.
+It addresses the pressing issue of hallucination in FMs, acknowledging its widespread impact across various domains.
+The paper underscores the importance of addressing this challenge by examining recent advancements in detection and mitigation techniques, given FMs’ indispensable role in critical tasks.
+Its primary contribution is presenting a structured taxonomy for classifying hallucination in FMs, spanning text, image, video, and audio domains.
+8 Limitation
 Previous survey papers primarily focused on hallucination in large language models and did not extensively cover hallucinations in vision, audio, and video modalities.
-In this survey paper, our aim is to provide a comprehensive overview of hallucinations across all modalities, considering that hallucinations can occur in any large foundation model.
-Despite our efforts to provide a comprehensive summary of recent advancements related to hallucination techniques in all foundational models, we acknowledge that we may miss some relevant work in the field.
-8 Future Directions
-Researchers are actively investigating hallucination mitigation techniques as the challenge of hallucination can be crucial in sensitive areas where generating fictional or incorrect content could have serious consequences Tonmoy et al.
-(2024)Rawte et al.
-(2023b).
-Here are the potential directions for addressing this critical issue of hallucination in these FMs:
-Data Resources: Recent studies have highlighted the efficacy of simple fine-tuning on carefully curated high-quality samples for reducing hallucinations, surpassing the impact of large-scale fine-tuning and reinforcement learning approaches.
-For knowledge-intensive domains, the development of entity-centered fine-tuned instructions that integrate structured knowledge from knowledge graphs shows promise in enhancing accuracy and relevance.
-Additionally, employing alignment techniques tailored to specific tasks or fields has proven effective in mitigating hallucinations.
-As research in this area progresses, more resources focused on improving alignment through task-specific or domain-adapted approaches are expected, further bolstering the reliability of language models in generating factual and trustworthy content.
-Automated Evaluation: Developing specialized evaluation metrics that consider factors such as factual accuracy and coherence can be useful for hallucination detection.
-Combining automated evaluation with human judgments via crowdsourcing can capture nuanced aspects challenging for automated systems alone.
-Additionally, adversarial testing methodologies are also being developed to expose AI systems to crafted inputs, aiding in identifying weaknesses and enhancing resilience against hallucination.
-Furthermore, fine-tuning FMs on datasets emphasizing fact-checking and accuracy offers another avenue to improve content reliability and reduce the occurrence of hallucinations.
-Improving Detection and Mitigation techniques: Mitigating hallucinations in FMs necessitates a multifaceted approach that leverages reasoning mechanisms, knowledge graph integration, specialized fact-checking models, bias mitigation techniques, and active learning methodologies.
-Emerging techniques like Chain of Thought (CoT) Wei et al.
-(2022) and Tree of Thought (ToT) Yao et al.
-(2024) bolster these models’ reasoning capabilities, potentially reducing hallucinations.
-Integrating knowledge graphs enhances understanding of factual information and concept relationships, aiding content generation and fact-checking.
-Specialized verification models cross-reference outputs against curated knowledge to identify inaccuracies, while bias detection and mitigation techniques promote fairness.
-Finally, ethical guidelines and regulatory frameworks governing the responsible use of curated knowledge in AI development mitigate risks and foster public trust, collectively improving the quality, accuracy, and trustworthiness of AI-generated content.
-Multimodal Hallucination: Addressing hallucinations in multimodal large Foundation models requires a comprehensive approach spanning data-centric initiatives, cross-modal alignment efforts, architectural innovations, standardized benchmarking, reframing of hallucination, and enhancing interpretability and trust.
-Data-centric techniques for robust data collection, augmentation, and calibration ensure diverse and high-quality training data.
-Cross-modal alignment focuses on aligning representations across modalities through sophisticated architectures.
-Model architectural advancements involve designing specialized models capable of handling complex linguistic and visual inputs effectively.
-Establishing unified metrics and standardized benchmarks enables accurate assessment of hallucination and reliable performance evaluations.
-Reframing hallucination as a feature explores its integration into downstream applications, optimizing for the human experience.
-Finally, developing techniques for interpreting model behavior, visualizing internals, and improving reliability assessment fosters trust in MLLMs.
-This multifaceted approach collectively addresses the critical hallucination challenge, paving the way for more reliable and trustworthy multimodal AI systems.
-9 Conclusion
-This survey paper systematically categorizes existing research on hallucination within FMs, providing comprehensive insights into critical facets, including detection, mitigation, tasks, datasets, and evaluation metrics.
-It addresses the pervasive impact of hallucination in FMs, acknowledging its impact across diverse domains.
-By examining recent advancements in detection and mitigation techniques, the paper underscores the importance of addressing this challenge, given FMs’ indispensable role in critical tasks.
-Its primary contribution lies in introducing a structured taxonomy for classifying hallucination in FMs, spanning text, image, video, and audio domains.
+This survey paper aims to provide a comprehensive overview of hallucinations across all modalities, considering that hallucinations can occur in any large foundation model.
+Despite our efforts to provide a comprehensive summary of recent advancements related to hallucination techniques in all foundational models, we acknowledge that we may miss some relevant work in the field and have covered the papers till May 2024.